### v1

# Quantization of Lie-Poisson algebra and Lie algebra solutions of mass-deformed type IIB matrix model

Abstract
A quantization of Lie-Poisson algebras is studied. The mass-deformed IIB matrix model admits classical solutions constructed from the basis of any semisimple Lie algebra. We consider the geometry described by the classical solutions of the Lie algebras in the limit where the mass vanishes and the matrix size is infinite. Lie-Poisson varieties are regarded as such geometric objects. We provide a quantization called “weak matrix regularization” of any Lie-Poisson algebra (linear Poisson algebra) on the algebraic variety defined by its Casimir polynomials. The Lie algebra that gives weak matrix regularization is not necessarily semisimple. Casimir polynomials correspond with Casimir operators of the Lie algebra by the quantization. This quantization is a generalization of the fuzzy sphere. In order to define the weak matrix regularization of the quotient space by the ideal generated by the Casimir polynomials, we take a construction method that fixes a reduced Gröbner basis of the ideal. The Gröbner basis determines remainders of polynomials. The operation of replacing this remainders with representation matrices of a Lie algebra roughly corresponds to a weak matrix regularization. As concrete examples, we construct weak matrix regularization for $\mathfrak{su}(2)$ and $\mathfrak{su}(3)$. In the case of $\mathfrak{su}(3)$, we not only construct weak matrix regularization for the quadratic Casimir polynomial, but also construct weak matrix regularization for the cubic Casimir polynomial.
[…]
1 Tokyo University of Science,
1-3 Kagurazaka, Shinjuku-ku, Tokyo, 162-8601, Japan
1 Introduction
In M-theory and string theory, it has been proposed for a long time that matrix models give their constructive formulation [19, 42]. Naturally, their physical quantities of classical solutions, etc. are given as matrices. (See for example [29, 80] and references therein.) If we assume that the universe we live in can be obtained as a set of matrices, then at least the universe we observe can be approximated by a smooth manifold, so we need a correspondence between matrices and smooth spacetime. In this paper, we will discuss quantization as a method for obtaining such a correspondence, and we will give a quantization of Poisson algebra varieties with Lie-Poisson structures.
The matrix model given as a constructive formulation of type IIB superstring theory is called the type IIB matrix model or the IKKT (Ishibashi, Kawai, Kitazawa, and Tsuchiya) matrix model [42]. In this paper, we will refer to this model as the IKKT matrix model. There are many studies that consider matrix solutions and eigenvalue distributions in the IKKT matrix model as spacetime. One representative approach is the use of numerical methods using computers to simulate eigenvalue distributions, with many studies discussing, for example, the creation of a $3+1$-dimensional spacetime [50]. Many results of numerical approaches have been summarized in [7], and more can be found by consulting the references therein. There are also many studies that view the effective theory of the IKKT matrix model as a field theory on a noncommutative space. The idea has its roots in earlier work, dating back to [36]. The observation of the connection between IKKT and noncommutative geometry was made in [29, 8]. The correspondence between finite-dimensional IKKT matrix models and gauge theories on noncommutative spaces was later revealed [4, 5, 6]. There are also several survey papers on this topic [80, 87]. As evidenced by the fact that the IKKT matrix model can be formulated as a zero-dimensional reduction of gauge theory, the model does not include a mass term, naively. However, it is known that the only classical solution for the finite-dimensional IKKT matrix model without a mass term is given as a set of simultaneously diagonalizable matrices. (See the appendix of [81].) The mass term in the IKKT matrix model was discussed in the new regularization of the model respecting the Lorentz symmetry in [17]. There is also research that suggests that the mass term itself exists effectively [56]. There have also been reports of attempts to derive gravity using the IKKT matrix model, which is a model that undergoes mass deformations that preserve supersymmetry [53, 54]. These deformations, however, are different from those that add a simple mass term. This paper discusses classical solutions of the IKKT matrix model with a mass term and the space they are expected to represent.
In those matrix models, a fuzzy space is an important concept that maps classical geometry to noncommutative matrix algebra. The fuzzy sphere is a typical example [61]. It provides a mapping from a set of polynomials on $S^{2}$ to the space of endomorphisms on an $N$-dimensional vector space, which is a matrix algebra generated by an irreducible representation of $\mathfrak{su}(2)$. This mapping provides an approximate correspondence between the Poisson brackets and commutators. Other fuzzy spaces as known fuzzy Riemann surfaces are constructed, due to the motivation of the membrane theory in [12, 15, 10, 9, 11, 74]. In a similar study with Toeplitz operators, [52] proved that a Poisson algebra on any Riemann surface is produced in the limit $N\to\infty$, and [26] extended it to compact Kähler manifolds. These methods of mapping classical geometry to matrix algebra are called matrix regularization. Recently, fuzzy spaces have been studied using quasi-coherent states, and it has been shown that it is possible to extract the classical structure without the $N\to\infty$ limit [74, 82]. This direction of research is developing in areas that are not limited to spaces with positive-valued measures. Chany, Lu and Stern extended the fuzzy sphere studied in Euclidean to Minkowski and showed that it is a solution to the Lorentzian IKKT matrix model [84]. Ho and Li showed that fuzzy $S^{2}$ and $S^{4}$ are candidates for the quantum geometry on the corresponding spheres in AdS/CFT correspondence [40]. [27] and subsequent papers by the same group report on fuzzy spaces in de Sitter and anti-de Sitter spaces.
Fuzzy spaces have not only been studied in the context of the IKKT matrix model or the BFSS matrix model, but also as a direction in noncommutative geometry. As already mentioned above, the history of fuzzy space begins with the proposal of the fuzzy sphere [41, 61]. We can find more details in the [41, 61, 10, 21], as well as in the references within them. Fuzzy torus is also constructed in a similar way. (See for example [20].) It is not possible to mention all examples, but it is possible to find summarized descriptions in, for example, Steinacker’s textbook [83]. In this paper we construct a method to obtain fuzzy spaces of Lie-Poisson algebras. As concrete examples, we also construct examples for $\mathfrak{su}(2)$ and $\mathfrak{su}(3)$. Closely related to the examples are, for example, fuzzy $CP^{2}$ [63, 38, 3, 22, 18, 37]. The Lie-Poisson algebras treated in this paper are not limited to those corresponding to compact groups, but it is also possible to consider Lie algebras for noncompact groups. The fuzzy spaces corresponding to noncompact groups are known, for example [39, 46, 47, 78, 79]. These various studies constructed fuzzy spaces somewhat ad hoc for individual manifolds. In contrast, the study by [71] construct general framework and mathematically precise statements by using $C^{*}$-metric, Gromov-Hausdorff-type distance. The purpose of constructing a fuzzy space of a coadjoint orbit of [71] is a slightly similar to the purpose of this paper, which is to quantize the space containing coadjoint orbit as a subspace. The method used in this paper is simpler and easier to compute. Various approaches to the correspondence between manifolds or algebraic varieties and the elements of matrix algebras or its subalgebras are expected be explored in the future, and this paper contributes to that effort.
Taking the classical(commutative) limit of a noncommutative manifold is often fraught with difficulty. For example, as a well-known phenomenon, matrix regularization generates the same matrix algebra whether a two-dimensional torus is transformed into a fuzzy torus or a two-dimensional sphere into a fuzzy sphere. (See [31, 14, 41, 61, 62].) Put another way, when reading a classical geometry (Poisson algebra) from a matrix algebra, the Poisson algebra obtained depends on what classical(commutative) limit is taken [26, 25, 28]. Sometimes, such the problem of determining the classical(commutative) limit is called an inverse problem. Various approaches have been taken to the inverse problem of how to extract geometric properties from matrix algebras [77, 24, 74, 44, 45, 16]. Therefore, a framework including any Poisson algebra and its quantizations is important in considering such issues to be investigated in a unified manner. In [72], a category of Poisson algebras and their quantized spaces is proposed, and a category-theoretic formulation of their classical limit is given. As an example, a framework for obtaining a Poisson algebra as the classical limit of any semisimple Lie algebra was discussed in it.
One purpose of this paper is to examine in detail the contents of Section 6 of that paper [72]. There, the quantization of Lie-Poisson algebras and their inverse problems, as well as derivations based on the principle of least action, were discussed. The definition of Lie-Poisson algebras in this paper is given as follows.
Definition 1.1.
Let $x=(x_{1},x_{2},\cdots,x_{d})$ be commutative variables. Suppose that $(\mathbb{C}[x],\cdot,\{~{},~{}\})$ is a Poisson algebra. If the Poisson bracket acts as a linearly i.e., there exist structure constants $f_{ij}^{k}$ such that $\{x_{i},~{}x_{j}\}=f_{ij}^{k}x_{k}$, we call $(\mathbb{C}[x],\cdot,\{~{},~{}\})$ Lie-Poisson algebra.
The Lie-Poisson structure is introduced in [59, 88]. The study of the quantization of Lie-Poisson algebras is given by Rieffel as a deformation quantization [69]. As a related study, deformation quantization of polynomial Poisson algebras via universal enveloping algebra (generalizing that of Lie-Poisson structures) is discussed in [67]. In [72], one of the authors discussed matrix regularization of Lie-Poisson algebras as an example of how the inverse problem of quantization can be treated in the category theoretical limit. However, detailed discussions were omitted, and the case of taking a quotient by a nontrivial ideal was not sufficiently addressed. In this paper, the details are given explicitly in general Lie algebras, and new concrete examples are constructed.
Here, a synopsis of this paper is presented.
At first, the equivalence of the mass-deformed IKKT matrix model and the matrix models whose classical solutions are given as semisimple Lie algebras is shown in Section 2. (However, only the Bosonic part is mentioned.) From the matrix model, it is shown that the mass-deformed IKKT matrix model has a classical solution constructed from a representation of any semisimple Lie algebra. This result was already derived by Arnlind and Hoppe in [13]. These facts make it an important topic to investigate the spaces that emerge as commutative limits of the solutions. When taking the commutative limit, we simultaneously take the limit in which the dimension of the representation of the Lie algebra tends to infinity.
In this paper, we consider Lie-Poisson algebras as such spaces. The algebra generated by a representation of Lie algebras is constructed by quantization of Lie-Poisson algebras. The space in which the Lie-Poisson algebra is defined is the variety described by $k$-th degree Casimir polynomials. The space includes a coadjoint orbit of the Lie algebra. The matrix regularization used in this paper is slightly different from the matrix regularization that is commonly used, and the conditions are weakened. Therefore, to distinguish between the two, the term “weak matrix regularization” is also used. The difference is that weak matrix regularization is defined algebraically without using an operator norm, and only the condition corresponding to quantization is retained. Details of the difference between matrix regularization and weak matrix regularization are discussed in Section 3.2. An equivalent class of a polynomial defined on the variety corresponds to a matrix by the weak matrix regularization. The Lie algebra that gives weak matrix regularization is not necessarily semisimple.
Here, we also mention an overview of how to construct the actual weak matrix regularization for a Lie-Poisson algebra ${\mathbb{C}}[x]/I$ given in this paper, where $I$ is an ideal generated by Casimir polynomials. Consider $[f(x)]\in{\mathbb{C}}[x]/I$. A reduced Gröbner basis $G$ of $I$ is introduced. Then for any $f(x)\in{\mathbb{C}}[x]$ $f(x)=r(x)+h(x)$ is uniquely determined, where $h(x)\in I$ and $r(x)\notin I$. Roughly speaking, the weak matrix regularization is achieved by replacing all the variables $x_{1},\cdots,x_{d}$ in this polynomial $r(x)$ with the corresponding elements of the basis of the representation matrix of the Lie algebra. This process is a generalization of the method of constructing the fuzzy sphere. In fact, as an example, we construct a fuzzy sphere by matrix regularization of $\mathfrak{su}(2)$ by the way of this paper. Furthermore, we consider an example of $\mathfrak{su}(3)$ and provide a matrix regularization of varieties determined not only by quadratic but also by cubic Casimir polynomials.
The organization of this paper is as follows. In Section 2, it is shown that any semisimple Lie algebra is a solution of the mass-deformed IKKT matrix model when the Killing form of the Lie algebra and the metric of the IKKT matrix model are same type. In Section3, we summarize review of [72] and some mathematical facts, the minimum necessary to be used in constructing the weak matrix regularization of Lie-Poisson algebras in this paper. Furthermore, the matrix regularization of varieties corresponding to the case where the ideal is trivial, i.e., $\{0\}$, is discussed. In Section 4, the weak matrix regularization of varieties that is determined by $k$-th degree Casimir polynomials is constructed. In Section 5, as examples, we construct weak matrix regularizations for the cases where the Lie algebra is $\mathfrak{su}(2)$ and $\mathfrak{su}(3)$. Section 6 provides a summary of this paper.
This paper basically uses Einstein summation convention, but in cases where it is difficult to understand, the summation symbol $\sum$ will be used as appropriate.
2 Lie algebras as solutions of IKKT matrix model
In this section, we discuss the relationship between mass-deformed IKKT matrix models and semisimple Lie algebras in order to clarify one of the motivations for this paper. Let us consider the Bosonic part of the IKKT matrix model with a mass regularization term. Using $N\times N$ Hermitian matrices $X^{N}_{\mu}~{}(\mu=1,\cdots,d,~{}N\in{\mathbb{N}})$ and a mass $\hbar(N)$, we consider the action
$S_{IKKT}(\hbar^{2}(N))[X]$ $=\mbox{\rm{tr}}\left(-\frac{1}{4}[X^{N}_{\mu},X^{N}_{\nu}]^{2}+\hbar^{2}(N)\frac{1}{2}X^{N\mu}X^{N}_{\mu}\right).$ (2.1)
Here we take contraction as
$[X^{N}_{\mu},X^{N}_{\nu}]^{2}=\eta^{\mu\rho}\eta^{\nu\tau}[X^{N}_{\mu},X^{N}_{\nu}][X^{N}_{\rho},X^{N}_{\tau}]=\sum_{\mu,\nu,\rho,\tau}\eta^{\mu\rho}\eta^{\nu\tau}[X^{N}_{\mu},X^{N}_{\nu}][X^{N}_{\rho},X^{N}_{\tau}].$
$\eta$ is a diagonal matrix and usually a Euclidean or a Minkowski metric. In this paper, however, $\eta$ is a more general case that also includes $(n,d-n)$-type Minkowski metric; $\eta={\rm diag}(1,\cdots,1,-1,\cdots,-1)$. The second term in (2.1) represents the deformation due to the mass term. It should really be called the “mass-deformed IKKT matrix model”, but in this paper we sometimes simply call it the IKKT matrix model. (As already mentioned in Section 1, it is known that there are various classical solutions when the model is deformed with a mass term. See for example [49, 51, 79].)
Remark.
A technical note should be made here. Although we stated that $X^{N}_{\mu}$ is a Hermitian matrix care must be taken when the matrix size $N$ is finite. Because of the problem of the lack of finite-dimensional Hermitian representations for noncompact groups, in the following, $X_{\mu}$ corresponding to the negative sign of the metric is represented by an anti-Hermitian matrix $iX_{\mu}$ as in the Wick rotation. In other words, since we consider a number of anti-Hermitian matrices corresponding to the negative eigenvalues, the model essentially corresponds to the Euclidean IKKT matrix model in finite dimensions. To make this explicit, we refer to it as the “Euclidean IKKT matrix model”. By doing Wick rotation, the following argument can be applied to any Killing metric with negative eigenvalues. Therefore, for simplicity, we will describe the case of positive-valued metrics without the imaginary unit.
In [72], classical limits of quantizations were discussed in the category “Quantum world” which contains the whole of classical spaces (Poisson algebras) and their quantized spaces. In it, a matrix model which treats the Lie algebras as quantum spaces was introduced. The matrix model was given as
$S_{N}(\hbar^{2}(N))[X]$ $=\mbox{\rm{tr}}\left(-\frac{1}{4}g^{\mu\rho}g^{\nu\tau}[X^{N}_{\mu},X^{N}_{\nu}][X^{N}_{\rho},X^{N}_{\tau}]+\hbar^{2}(N)\frac{1}{2}g^{\mu\nu}X^{N}_{\mu}X^{N}_{\nu}\right).$ (2.2)
Here $g^{\mu\nu}\in{\mathbb{R}}$ represents the component of a real symmetric nondegenerate matrix. In the following, $g$ is referred to as a metric for simplicity. (There have been similar studies in the past that have the same equation of motion [43, 48], but the action themselves are different.) It is possible to show that the model (2.2) is actually equivalent to the IKKT matrix model (2.1). We shall first look at derivation of this fact.
When considering ordinary Riemannian geometry, there exist a local coordinate transformation that can transform any nondegenerate symmetric positive matrix into the Euclidean metric as $g_{\mu\nu}dx^{\mu}dx^{\nu}=\eta_{\mu\nu}dx^{\prime\mu}dx^{\prime\nu}$. In the same way now, we diagonalize the metric $g$ and perform the variable transformation as follows.
$\sum_{\mu,\nu}{O^{-1}}_{\tau\mu}g^{\mu\nu}{O_{\nu\rho}}=\lambda_{\tau}\eta^{\tau\rho},\qquad\sum_{\nu}\sqrt{\lambda_{\tau}}{O^{-1}}_{\tau\nu}X^{N}_{\nu}=\sum_{\nu}\sqrt{\lambda_{\tau}}{O}_{\nu\tau}X^{N}_{\nu}=:X^{N\prime}_{\tau}.$ (2.3)
Einstein’s contraction notation is not used here. In other words, no contraction is taken with respect to the subscript $\tau$. Here $O$ is an orthogonal matrix to diagonalize $(g^{\mu\nu})$, and $\lambda_{\tau}\eta^{\tau\tau}~{}(\tau=1,\cdots,d)$ are eigenvalues of the matrix $(g^{\mu\nu})$. Note that $X^{N\prime}_{\tau}$ remain a Hermitian matrix. Using this variable changing,
$g^{\mu\nu}X^{N}_{\mu}X^{N}_{\nu}=\eta^{\mu\nu}X^{N\prime}_{\mu}X^{N\prime}_{\nu}$ (2.4)
and
$g^{\mu\rho}g^{\nu\tau}[X^{N}_{\mu},X^{N}_{\nu}][X^{N}_{\rho},X^{N}_{\tau}]=\eta^{\mu\rho}\eta^{\nu\tau}[X^{N\prime}_{\mu},X^{N\prime}_{\nu}][X^{N\prime}_{\rho},X^{N\prime}_{\tau}]$ (2.5)
are easily obtained. Then we get
$S_{IKKT}(\hbar^{2}(N))[X^{\prime}]=S_{N}(\hbar^{2}(N))[X].$ (2.6)
Remark.
The path integral measure of the Bosonic part of the IKKT matrix model is given by
${\mathcal{D}}X:=\prod_{\mu}\left(\prod_{i=1}^{N}d(X^{N}_{\mu})_{ii}\right)\left(\prod_{i>j}d(X^{NRe}_{\mu})_{ij}d(X^{NIm}_{\mu})_{ij}\right).$
Here we calculate the Jacobian that appears in the above variable changing. From (2.3),
$d(X^{N}_{\mu})_{ij}=\sum_{\tau}\frac{1}{\sqrt{\lambda_{\tau}}}{O_{\mu\tau}}d(X^{N\prime}_{\tau})_{ij}.$
Note that
$\det\left(\frac{1}{\sqrt{\lambda_{\nu}}}{O_{\mu\nu}}\right)=\prod_{\tau=1}^{d}\frac{1}{\sqrt{\lambda_{\tau}}}\det\left({O_{\mu\nu}}\right)=\sqrt{\det(g_{\mu\nu})},$
where $(g_{\mu\nu})$ is the inverse matrix of $(g^{\mu\nu})$. Then we obtain
${\mathcal{D}}X=(\det(g_{\mu\nu}))^{\frac{N^{2}}{2}}{\mathcal{D}}X^{\prime}.$ (2.7)
So, if we define the Bosonic part partition functions
$Z_{IKKT}:=\int{\mathcal{D}}Xe^{-S_{IKKT}},\qquad Z_{Lie}:=\int{\mathcal{D}}Xe^{-S_{N}},$
the relation between them is the following;
Proposition 2.1.
The relation between the two Bosonic part partition functions given by the actions (2.1) and (2.2) is as follows.
$Z_{IKKT}=(\det(g_{\mu\nu}))^{\frac{N^{2}}{2}}Z_{Lie}.$ (2.8)
Next, we consider the case where the metric $g$ is given by the Killing metric of Lie algebra. The Killing metric is expressed as $g_{\mu\nu}=-f_{\mu\rho}^{\tau}f_{\nu\tau}^{\rho}$, where $f_{\mu\rho}^{\tau}$ are structure constants of some Lie algebra $\mathfrak{g}$. Since the existence of this non-degenerate metric is equivalent to the semisimplicity of the Lie algebra, we require that the Lie algebra $\mathfrak{g}$ is semisimple in the following discussion of the Killing form as a metric.
As already shown in [72], the equation of motion of (2.2) ,
$[X^{N\mu},[X^{N}_{\mu},X^{N}_{\nu}]]=-\hbar^{2}(N)X^{N}_{\nu},$ (2.9)
has the solution such that
$[X^{N}_{\mu},X^{N}_{\nu}]=\hbar(N)f_{\mu\nu}^{\rho}X^{N}_{\rho},$ (2.10)
with an orthogonal $\mathfrak{g}$ basis $\{X^{N}_{\mu}\}$ satisfying ${\rm tr}X^{N\mu}X_{\nu}^{N}=g^{\mu\tau}{\rm tr}X^{N}_{\tau}X_{\nu}^{N}=c{\delta^{\mu}}_{\nu}$, where $c$ is a constant. In fact, we substitute the commutation relation (2.10) for the left side of (2.9).
$[X^{N\mu},[X^{N}_{\mu},X^{N}_{\nu}]]$ $=\hbar(N)f_{\mu\nu}^{\rho}g^{\mu\tau}[X^{N}_{\tau},X^{N}_{\rho}]=\hbar^{2}(N)f_{\mu\nu}^{\rho}g^{\mu\tau}f_{\tau\rho}^{\sigma}X^{N}_{\sigma}=\hbar^{2}(N)f_{\mu\nu}^{\rho}g^{\mu\tau}g^{\eta\sigma}f_{\tau\rho\eta}X^{N}_{\sigma}$
$=\hbar^{2}(N)f_{\nu\mu}^{\rho}g^{\mu\tau}g^{\eta\sigma}f_{\eta\rho\tau}X^{N}_{\sigma}=-\hbar^{2}(N){\delta_{\nu}}^{\sigma}X^{N}_{\sigma}=-\hbar^{2}(N)X^{N}_{\nu}.$
Here we used the property that $f_{\mu\nu\rho}:=f_{\mu\nu}^{\tau}g_{\tau\rho}$ is totally anti-symmetric in the three indices when we chose the basis ${\rm tr}X^{N\mu}X_{\nu}^{N}=c{\delta^{\mu}}_{\nu}$. Therefore, we found that such set of generators $\{X^{N}_{\mu}\}$ which consist a representation of $\mathfrak{g}$ is a solution of the equation of motion of the action $S_{N}$.
Summarizing the above considerations, the following theorem is obtained.
Theorem 2.2.
Let $\mathfrak{g}$ be a semisimple Lie algebra. The mass-deformed Euclidean IKKT matrix model has a classical solution constructed from a representation of any $\mathfrak{g}$. The solution $X^{\prime}$ is constructed by using a $\mathfrak{g}$ basis $X$;
$X^{N\prime}_{\tau}=\sum_{\nu}\sqrt{\lambda_{\tau}}{O^{-1}}_{\tau\nu}X^{N}_{\nu},$ (2.11)
where $O$ is an orthogonal matrix such that $\sum_{\mu,\nu}{O^{-1}}_{\tau\mu}g^{\mu\nu}{O_{\nu\rho}}=\lambda_{\tau}\eta^{\tau\rho}$, and the basis $X$ satisfies
$[X^{N}_{\mu},X^{N}_{\nu}]=\hbar(N)f_{\mu\nu}^{\rho}X^{N}_{\rho},\quad{\rm tr}X^{N\mu}X_{\nu}^{N}=c{\delta^{\mu}}_{\nu},\quad g_{\mu\nu}=-f_{\mu\rho}^{\tau}f_{\nu\tau}^{\rho}.$
The statement of this theorem is not new. Although it is phrased differently, essentially the same result was derived by Arnlind and Hoppe in [13].
As an aside, it should also be noted that structure constants are also change as a result of variable transformation. To be more specific, if we define the structure constants $f_{\mu\nu}^{\prime\rho}$ as $[X^{N\prime}_{\mu},X^{N\prime}_{\nu}]=\hbar(N)f_{\mu\nu}^{\prime\rho}X^{N\prime}_{\rho}$, then this structure constants are given by
$f_{\mu\nu}^{\prime\rho}=\sqrt{\frac{\lambda_{\mu}\lambda_{\nu}}{\lambda_{\rho}}}\sum_{\alpha,\beta,\sigma}O_{\alpha\mu}O_{\beta\nu}{O^{-1}}_{\rho\sigma}f^{\sigma}_{\alpha\beta}.$ (2.12)
As described in this section, we have seen that the Lie algebra gives the solution of the IKKT matrix model with mass deformation. Therefore, it is important to investigate classical objects for which quantization yields a Lie algebra or an algebra generated by a representation of that Lie algebra. In the following, we consider this issue.
3 Quantization and preparations
In the previous section, we saw that any basis of an arbitrary semisimple Lie algebra corresponds to a classical solution of the mass-deformed IKKT matrix model. The algebra generated by the representation of the Lie algebra corresponds to a noncommutative spacetime. It is expected that Lie-Poisson algebras appear as the spaces in commutative limits. In Sections 3 and beyond, we consider the mechanism that the noncommutative spacetime is obtained by quantization of a Lie-Poisson algebra. If we only consider the classical solutions of the mass-deformed IKKT matrix model, then the Lie algebra must satisfy the condition of being “semisimple”, but the following discussion is valid for general Lie algebras.
In order to construct the quantization of Lie-Poisson algebras, we review some of the contents of [72, 73] below, and prepare for the discussion in Sections 4 and beyond.
3.1 Quantization
The following definition of quantization is used in this paper.
Definition 3.1 (Quantization map $Q$ [72] ).
Let ${A}$ be a Poisson algebra over a commutative ring $R$ over ${\mathbb{C}}$, and let $T_{i}$ be an $R$-module that is given by a subset of some $R$-algebra $(M,*_{M})$. If an $R$-module homomorphism (linear map) $t_{Ai}:{A}\to T_{i}$ equips a constant $\hbar(t_{Ai})\in{\mathbb{C}}-\{0\}$ and satisfies
$[t_{Ai}(f),t_{Ai}(g)]_{M}=\sqrt{-1}\hbar(t_{Ai})~{}t_{Ai}(\{f,g\})+\tilde{O}(\hbar^{1+\epsilon}(t_{Ai}))\quad(\epsilon>0)$ (3.1)
for arbitrary $f,g\in{A}$, where $[a,b]_{M}:=a*_{M}b-b*_{M}a$, we call $t_{Ai}$ a quantization map or simply a quantization. $\tilde{O}$ is defined in the Appendix A. We denote the set of all quantization maps by $Q$.
In this paper, we use ${\mathbb{C}}$ or ${\mathbb{C}}[\hbar]$ as $R$ below. One may wonder why the symbol $\tilde{O}$, which differs from the usual Landau symbol, is introduced in this definition of quantization. The reason why $\tilde{O}$ is used is that a norm is not introduced in $T_{i}$, so the norm of the scalar ${\mathbb{C}}$ is used to determine the limit. See also Appendix A and Appendix B for a detailed explanation.
In [72], the author defined the quantization as a part of the category of Poisson algebras and their quantizations, but in this paper, the discussion using categories is not necessary, so that part has been omitted.
The above definition of quantization includes many kinds of quantizations. For example, matrix regularization [41, 14], fuzzy spaces [61], and Berezin-Toeplitz quantization [26, 25, 75] which have original ideas of the matrix regularization, satisfy it. In addition, the strict deformation quantization introduced by Rieffel [68, 70, 57], the prequantization [23, 58, 86], and Poisson enveloping algebras [64, 65, 85, 89] are also in $Q$. (In [34, 35], we can see organized discussions about these quantization maps. The conditions for $Q$ are a part of the definition of pre-$\mathscr{Q}$ in [34, 35].)
3.2 Matrix regularization and fuzzy spaces for Lie-Poisson algebras
The matrix regularization for Lie-Poisson algebras is introduced in [72]. We review and refine it as “weak matrix regularization”. Furthermore, we discuss approximate algebra homomorphism between Lie-Poisson algebras and algebras generated by Lie algebras in this subsection.
Let $\mathfrak{g}$ be a finite dimensional Lie algebra. Let $e=\{e_{1},e_{2},\cdots,e_{d}\}$ be a fixed basis of $\mathfrak{g}$ satisfying commutation relations $[e_{i},e_{j}]=f_{ij}^{k}e_{k}$, where $f_{ij}^{k}$ are structure constants of $\mathfrak{g}$. For this Lie algebra $\mathfrak{g}$ we introduce an irreducible representation $\rho^{\mu}:\mathfrak{g}\rightarrow gl(V^{\mu})$ and a number $\hbar({\mu})\neq 0$. Here $V^{\mu}$ is a finite dimensional vector space chosen as appropriate. We denote the corresponding basis of $e$ by
$e^{(\mu)}=\{\hbar(\mu)\rho^{\mu}(e_{1}),\hbar(\mu)\rho^{\mu}(e_{2}),\cdots,\hbar(\mu)\rho^{\mu}(e_{d})\}=\{e_{1}^{(\mu)},e_{2}^{(\mu)},\cdots,e_{d}^{(\mu)}\}.$ (3.2)
Then they satisfy
$[e_{i}^{(\mu)},e_{j}^{(\mu)}]=\hbar(\mu)f_{ij}^{k}e_{k}^{(\mu)}.$ (3.3)
The Lie algebra $\rho^{\mu}(\mathfrak{g})$ is constructed by this basis.
Next, we introduce a Poisson algebra corresponding to this Lie algebra. There is a well-known way known as Kirillov-Kostant Poisson bracket, that is the way constructing a Poisson algebra. (See for example [41, 55, 60, 76, 2].) Let $x=(x_{1},x_{2},\cdots,x_{d})$ be commutative variables. We consider a Lie-Poisson algebra $(\mathbb{C}[x],\cdot,\{~{},~{}\})$ by
$\{x_{i},~{}x_{j}\}:=f_{ij}^{k}x_{k},$ (3.4)
where $f_{ij}^{k}\in{\mathbb{C}}$ are structure constants. Concretely, this Poisson bracket is realized by
$\{f,~{}g\}:=f\omega g:=f\overleftarrow{\partial}_{i}\omega_{ij}\overrightarrow{\partial}_{j}g:=({\partial}_{i}f)\omega_{ij}({\partial}_{j}g),$ (3.5)
where $\partial_{i}=\frac{\partial}{\partial x_{i}}$ and $\omega_{ij}=f_{ij}^{k}x_{k}$. It is easy to verify that (3.5) satisfies the Leibniz’s rule and the Jacobi identity. We denote the Poisson algebra $(\mathbb{C}[x],\cdot,\{~{},~{}\})$ by $A_{\mathfrak{g}}$. We define degree of a monomial $x^{\alpha}=(x_{1})^{\alpha_{1}}(x_{2})^{\alpha_{2}}\cdots(x_{d})^{\alpha_{d}}$ by $\deg x^{\alpha}:=|\alpha|:=\sum_{i=1}^{d}\alpha_{i}$. For the polynomial $f(x)=\sum_{\alpha}a_{\alpha}x^{\alpha}$, where $a_{\alpha}\in\mathbb{C}$, $\deg f(x)$ is defined by $\max_{a_{\alpha}\neq 0}\{\deg x^{\alpha}\}$. For multi-index, the notation $x^{I}=x_{i_{1}}x_{i_{2}}\cdots x_{i_{m}}$ is also often used below, where $\deg x^{I}=m=:|I|$.
Next, let us construct quantization maps from the Lie-Poisson algebra $A_{\mathfrak{g}}$ to $T_{\mu}$ that is $\langle e^{(\mu)}\rangle$. We denote the $R$-algebra generated by $e^{(\mu)}$ by $\langle e^{(\mu)}\rangle$, here. (If $\mathfrak{g}$ is a semisimple Lie algebra, then $T_{\mu}$ is given by $End(V^{\mu})=gl(V^{\mu})$.) In [72], $T_{\mu}$ is regarded as a vector space that is $\langle e^{(\mu)}\rangle$ forgetting multiplication structure. However, we do not discuss the category $QW$ in this paper, so there is no problem treating $T_{\mu}$ as an algebra. We choose a basis of $T_{\mu}:=\langle e^{(\mu)}\rangle$, $E_{1},E_{2},\dots,E_{D}$, as polynomials of $e^{(\mu)}$. Any polynomial of $e^{(\mu)}$ can be rewritten by $\hbar$ polynomial in $\langle\rho^{\mu}(e)\rangle[\hbar]$. So, a degree ${\rm deg}$ of any polynomial of $e^{(\mu)}$ can be defined by $\hbar$’s degree. Using $E^{i}_{j_{1},\cdots,j_{k}}\in{\mathbb{C}}$, $E_{i}~{}(i=1,\cdots,D)$ are expressed as
$E_{i}=\sum_{k}E^{i}_{j_{1},\cdots,j_{k}}e^{(\mu)}_{j_{1}}\cdots e^{(\mu)}_{j_{k}}=\sum_{k}\hbar^{k}(\mu)E^{i}_{j_{1},\cdots,j_{k}}\rho^{\mu}(e_{j_{1}})\cdots\rho^{\mu}(e_{j_{k}}),$
where $E^{i}_{j_{1},\cdots,j_{k}}$ is independent of $\hbar$. Note that Einstein summation convention is used for each $j_{l}$. For each $E_{i}$, such expression given by $e^{(\mu)}$ is not unique in general. We chose a expression that minimizes $\max_{E^{i}_{j_{1},\cdots,j_{k}}\!\!\!\!\!\!\!\!\!\!\neq 0}\{k\}$. Then there exists the degree of $\hbar$ of $E_{i}$ i.e., $\deg E_{i}:=\max_{E^{i}_{j_{1},\cdots,j_{k}}\!\!\!\!\!\!\!\!\!\!\neq 0}\{k\}$. The highest degree of $\{E_{1},E_{2},\dots,E_{D}\}$ is denoted by $n_{\mu}$, i.e., $n_{\mu}=\max\{{\deg}E_{1},\cdots,{\rm deg}E_{D}\}$. $n_{\mu}$ does not depend on the choice of $\{E_{1},E_{2},\dots,E_{D}\}$, because another $E^{\prime}_{1},E^{\prime}_{2},\dots,E^{\prime}_{D}$ can also be described as linear combinations of the original $E_{1},E_{2},\dots,E_{D}$.
In considering matrix regularization, the quantization of the target space by a finite matrix algebra (and its subalgebras), the following properties are characteristic.
Lemma 3.2.
Let the vector $E_{i_{1}}E_{i_{2}}\cdots E_{i_{k}}$ be represented by a linear combination of $\{E_{1},E_{2},\dots,E_{D}\}$ with each component $c_{i}$.
$E_{i_{1}}E_{i_{2}}\cdots E_{i_{k}}=\sum_{j}c_{j}E_{j}.$
If $\sum_{l=1}^{k}\deg E_{i_{l}}>n_{\mu}$, then $c_{i}$ is a polynomial consisting of terms of degree $1$ or higher in $\hbar(\mu)$. As a similar claim, for $k>n_{\mu}$ and
$e^{(\mu)}_{i_{1}}\cdots e^{(\mu)}_{i_{k}}=\sum_{j}c_{j}E_{j},$
then $c_{i}$ is a polynomial consisting of terms of degree $k-n_{\mu}$ or higher in $\hbar(\mu)$.
Proof.
It follows from the fact that the order of $\hbar$ of $E_{i_{1}}E_{i_{2}}\cdots E_{i_{k}}$ is greater than $n_{\mu}$, but the maximum order of $E_{i}$ is $n_{\mu}$. ∎
Definition 3.3.
We define linear function $q_{\mu}:A_{\mathfrak{g}}\rightarrow T_{\mu}$ by
$\sum_{I}f_{I}x^{I}:=\sum_{k}f_{i_{1},\cdots,i_{k}}x_{i_{1}}\cdots x_{i_{k}}\mapsto\sum_{I}f_{I}e^{(\mu)}_{(I)}=\sum_{k}^{n_{\mu}}f_{i_{1},\cdots,i_{k}}e^{(\mu)}_{(i_{1},\cdots,i_{k})},$ (3.6)
where $f_{i_{1},\cdots,i_{k}}\in\mathbb{C}$ is completely symmetric,
$e^{(\mu)}_{(I)}:=e^{(\mu)}_{(i_{1},\cdots,i_{k})}:=\frac{1}{k!}\sum_{\sigma\in Sym(k)}e^{(\mu)}_{i_{\sigma(1)}}\cdots e^{(\mu)}_{i_{\sigma(k)}},$
and we require that the multiplicative identity of $A_{\mathfrak{g}}$ maps to the unit matrix in $T_{\mu}$ i.e., $q_{\mu}(1)=Id\in gl(V^{\mu})$.
Remark.
$n_{\mu}$ is introduced to give a boundary that determines the kernel of $q_{\mu}$. Although it is defined by $n_{\mu}=\max\{{\rm deg}E_{1},\cdots,{\rm deg}E_{D}\}$ here, there is no particular reason why it has to be this way. The important points are that $n_{\mu}$ is defined as a sequence that increases with the dimension of $V^{\mu}$, and that Lemma 3.2 is satisfied. But the details are not essential. For example, it is also possible to define $n_{\mu}$ by using multi degree as follows. We denote multi degree of $E_{i}$ by ${\rm multdeg}E_{i}$. (In Appendix C, we can see a definition of multi degree.) $n_{\mu}$ can be defined as the highest multi degree of $\{E_{1},E_{2},\dots,E_{D}\}$ i.e. $n_{\mu}=\max\{{\rm multdeg}E_{1},\cdots,{\rm multdeg}E_{D}\}$. It is possible to reduce the number of more degenerate elements of $q_{\mu}$. However, this definition has not been chosen for simplicity. Therefore, the notation $\sum_{k}^{n_{\mu}}$ used here and in the following should be more accurately interpreted as $\sum_{(i_{1},\cdots,i_{k})\in Dom}$, where
$Dom=\{(i_{1},\cdots,i_{k})~{}|~{}x_{i_{1}}\cdots x_{i_{k}}\in A_{\mathfrak{g}}\backslash\ker q_{\mu}\}.$ (3.7)
By definition, this quantization $q_{\mu}$ satisfies the following.
Proposition 3.4.
Let $q_{\mu}:A_{\mathfrak{g}}\rightarrow T_{\mu}=\langle e^{(\mu)}\rangle$ be a linear function of Definition 3.3. Then it satisfies
$[q_{\mu}(f),q_{\mu}(g)]=\hbar({\mu})q_{\mu}(\{f,g\})+\tilde{O}(\hbar^{2}({\mu}))$
for $\forall f,g\in A_{\mathfrak{g}}$. In other words, $q_{\mu}\in Q$.
The proof is given in [72], however, there is no need to refer to it, as we will give a more detailed discussion soon in a slightly different framework of “weak matrix regularization”. Since $q_{\mu}$ maps $n$-degree polynomials to $n$-degree quantities in $T_{\mu}$, this proposition looks almost a trivial assertion. In other words, there is the inability to distinguish between $\hbar$, which represents noncommutativity, and $\hbar$, which originates from the degree of the polynomial. This makes it insufficient to use the degree of $\hbar$ as a meaningful measure of noncommutativity. The quantization from a Lie-Poisson algebra to a matrix algebra or a subset of a matrix algebra, such as $q_{\mu}\in Q$, which is included in the quantization $Q$ is regarded as a matrix regularization in this paper. Since its construction is a certain generalization of the matrix regularization of Madore[61] or de Wit-Hoppe-Nicolai [31], this quantization is regarded as a matrix regularization. The target of matrix regularization is called a Fuzzy space. There is not necessarily a consensus on a single definition of “matrix regularization”. Commonly used definition of matrix regularization is given in Appendix B. The definition used in this paper is less restrictive than the one in Appendix B. We define weak matrix regularization here.
Definition 3.5 (Weak matrix regularization).
Consider a Poisson algebra $A$ and a sequence of subalgebras $B^{m}\subset End(V^{m})~{}(m=1,2,\cdots)$ where $V^{m}$ is a fnite dimensional vector space and $\lim_{m\rightarrow\infty}\dim V^{m}=\infty$. Let $x=(x_{i},\cdots,x_{d})$ be a generator of $A$. Let $q_{m}:A\to B^{m}~{}(m=1,2,\cdots)$ be a sequence of quantizations such that $\hbar(m):=\hbar(q_{m})$ tends to zero as the dimension of $V^{m}$ tends to infinity. For each $q_{m}$ and $\forall f,g\in A$, when there exists some
$P_{m}=\sum_{k}a^{m}_{i_{1},\cdots,i_{k}}q_{m}(x_{i_{1}})\cdots q_{m}(x_{i_{k}})\in B^{m},$
where $a^{m}_{i_{1},\cdots,i_{k}}\in{\mathbb{C}}$ does not depend on $\hbar(m)$ and $P_{m}$ is $\tilde{O}(\hbar^{k}(m))~{}(k\geq 0)$ satisfying
$[q_{m}(f),q_{m}(g)]=\hbar(m)q_{m}(\{f,g\})+\hbar^{2}({m})P_{m},$ (3.8)
we call $\{q_{m}\}$ a weak matrix regularization. We also simply say that $q_{m}$ is a weak matrix regularization.
Theorem 3.6.
Let $q_{\mu}$ be a quantization defined by Definition 3.3. Suppose that $\{\hbar(\mu)\}$ is a sequence such that $\hbar(\mu)\rightarrow 0$ as $\dim V^{\mu}\rightarrow\infty$ Then $q_{\mu}$ is a weak matrix regularization.
Proof.
If we can show that for $\forall f,g\in A_{\mathfrak{g}}$ there exist $P=\sum_{i}^{D}c_{i}(\hbar(\mu))E_{i}\in T_{\mu}$ with $c_{i}(\hbar)\in{\mathbb{C}}[\hbar(\mu)]$ such that
$[q_{\mu}(f),q_{\mu}(g)]=\hbar({\mu})q_{\mu}(\{f,g\})+\hbar^{2}({\mu})P,$ (3.9)
then $q_{\mu}$ is a matrix regularization. By Definition 3.3, $q_{\mu}(x^{I_{k}})=e^{(\mu)}_{(I_{k})}=\frac{1}{k!}\sum_{\sigma\in Sym(k)}e^{(\mu)}_{i_{\sigma(1)}}\cdots e^{(\mu)}_{i_{\sigma(k)}}$ for $|I_{k}|\leq n_{\mu}$, and $q_{\mu}(x^{I_{k}})=0$ for $|I_{k}|>n_{\mu}$.
When $|I_{k}|+|J_{m}|\leq n_{\mu}+1$, degree of $\{x^{(I_{k})},x^{(J_{m})}\}$ is $|I_{k}|+|J_{m}|-1$. So, $[e^{(\mu)}_{(I_{k})},e^{(\mu)}_{(J_{m})}]=\hbar(\mu)q_{\mu}(\{x^{(I_{k})},x^{(J_{m})}\})+\hbar^{2}(\mu)P_{1}$. Here we use $P_{k}=\sum_{i}^{D}c^{k}_{i}(\hbar(\mu))E_{i}\in T_{\mu}$, that is, each coefficient of base $E_{i}$ is a polynomial of $\hbar(\mu)$. When $|I_{k}|+|J_{m}|>n_{\mu}+1$, $q_{\mu}(\{x^{(I_{k})},x^{(J_{m})}\})=0$. $[e^{(\mu)}_{(I_{k})},e^{(\mu)}_{(J_{m})}]=\hbar^{2}(\mu)P_{2}$, since the commutator makes $\hbar(\mu)$ and the other $\hbar(\mu)$ arise from Lemma 3.2. Then, for $f=\sum_{I_{k}}f_{I_{k}}x^{I_{k}}$ and $g=\sum_{J_{m}}g_{J_{m}}x^{J_{m}}$,
$[q_{\mu}(f),q_{\mu}(g)]$ $=\sum_{|I_{k}|,|J_{m}|\leq n_{\mu}}f_{I_{k}}g_{J_{m}}[e^{(\mu)}_{(I_{k})},e^{(\mu)}_{(J_{m})}]$
$=\sum_{|I_{k}|+|J_{m}|\leq n_{\mu}+1}f_{I_{k}}g_{J_{m}}[e^{(\mu)}_{(I_{k})},e^{(\mu)}_{(J_{m})}]+\hbar^{2}(\mu)P_{3}$
$=\hbar(\mu)q_{\mu}(\{f,g\})+\hbar^{2}(\mu)P_{4}.$ (3.10)
$q_{\mu}(\{f,g\})$ is symmetrized, but reordered in the same order as $[e^{(\mu)}_{(I_{k})},e^{(\mu)}_{(J_{m})}]$ in the second line by using the commutation relation of the Lie algebra, and all the extra terms from the reordering are included in the second term. ∎
In the definition of weak matrix regularization, the restrictions on correspondence with the classical(commutative) limit are relaxed. How to choose the limit that brings $\hbar$ close to zero, or so-called classical(commutative) limit, is a delicate matter. Indeed, these classical limits lead to various classical Poisson algebras [28, 72]. Therefore, some restrictions on the classical limit are not be discussed in this paper. For example, quantization maps are not equipped with algebra homomorphism, but it is required to be homomorphic in the limit where $\hbar$ approaches zero in Appendix B. This paper does not focus on a rigorous proof of algebra homomorphism, but we discuss this issue briefly in the remainder of this section. We need a little preparation. First, let us introduce an enveloping algebra with variable $\hbar$.
From a set $X=\{X_{1},\cdots,X_{d}\}$ we make free monoid $\langle X\rangle=\{1\}\cup\{X_{i_{1}}X_{i_{2}}\cdots X_{i_{n}}~{}|~{}n\in\mathbb{N},X_{i_{j}}\in X\}$ and free ${\mathbb{C}}$-algebra
${\mathbb{C}}\langle X\rangle=\left\{\sum_{n}\sum_{{i_{1}},{i_{2}},\cdots,{i_{n}}}a_{{i_{1}},{i_{2}},\cdots,{i_{n}}}X_{i_{1}}X_{i_{2}}\cdots X_{i_{n}}\right\}.$
The usual enveloping algebra of $\mathfrak{g}$, $\mathcal{U}_{\mathfrak{g}}$, is an algebra of all polynomials of $X_{1},\cdots,X_{d}$ with relations $X_{i}X_{j}-X_{j}X_{i}\sim[X_{i}~{},~{}X_{j}]:=f_{ij}^{k}X_{k}$. This is a canonical definition of enveloping algebra. We introduce a slight changed algebra. Let $\mathcal{U}_{\mathfrak{g}}[\hbar]$ be an algebra $R\langle X\rangle$ over $R={\mathbb{C}}[\hbar]$ divided by a two-side ideal $I$ generated by
$X_{i}X_{j}-X_{j}X_{i}-[X_{i}~{},~{}X_{j}]:=X_{i}X_{j}-X_{j}X_{i}-\hbar f_{ij}^{k}X_{k}.$ (3.11)
$I:=\left\{\sum_{i,j,k,l}a_{k}(X_{i}X_{j}-X_{j}X_{i}-[X_{i}~{},~{}X_{j}])b_{l}~{}|~{}a_{k},b_{l}\in R\langle X\rangle\right\}.$ (3.12)
Note that $\hbar$ is introduced in the relation. Let us introduce the following.
$\mathcal{U}_{\mathfrak{g}}[\hbar]:=R\langle X\rangle/I.$
This enveloping algebra $\mathcal{U}_{\mathfrak{g}}[\hbar]$ allows for discussion of the rest of this section and the next section.
Every $E_{i}$, which is the basis of $T_{\mu}$, has an expression of
$E_{i}=\sum_{k}^{n_{\mu}}E^{(i)}_{j_{1},\cdots,j_{k}}e^{(\mu)}_{(j_{1},\cdots,j_{k})},$ (3.13)
where $E^{(i)}_{i_{1},\cdots,i_{k}}\in{\mathbb{C}}$ is completely symmetric coefficient. (In this paper, Einstein summation convention is used, so the above expression means $E_{i}=\sum_{k}^{n_{\mu}}\sum_{j_{1},\cdots,j_{k}}E^{(i)}_{j_{1},\cdots,j_{k}}e^{(\mu)}_{(j_{1},\cdots,j_{k})}.$) This is follows from the following Lemma, which is essentially equivalent to a part of Poincaré - Birkhoff - Witt (PBW) theorem.
Lemma 3.7.
Let $\mathcal{U}_{\mathfrak{g}}[\hbar]$ be the enveloping algebra of Lie algebra $\mathfrak{g}$ defined above. Then $\forall X_{i_{1}}X_{i_{2}}\cdots X_{i_{k}}\in\mathcal{U}_{\mathfrak{g}}[\hbar]$ can uniquely be written as
$X_{i_{1}}X_{i_{2}}\cdots X_{i_{k}}=X_{(i_{1},\cdots,i_{k})}+\sum_{l=0}^{k-1}a_{j_{1},\cdots,j_{l}}X_{(j_{1},\cdots,j_{l})},$ (3.14)
where $a_{j_{1},\cdots,j_{l}}\in\mathbb{C}$ is completely symmetric, and
$X_{(i_{1},\cdots,i_{l})}:=\frac{1}{l!}\sum_{\sigma\in Sym(l)}X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(l)}}.$
Proof.
We shall show by mathematical induction. $k=1$ is trivial. Assume that it is valid up to the $l$-th. It is sufficient to show that it is true for $X_{(i_{1},\cdots,i_{l})}X_{i_{l+1}}$.
$X_{(i_{1},\cdots,i_{l})}X_{i_{l+1}}=$ $\frac{1}{l!}\sum_{\sigma\in Sym(l)}X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(l)}}X_{i_{l+1}}$
$=$ $\frac{1}{l+1}\frac{1}{l!}\sum_{\sigma\in Sym(l)}\sum_{k=1}^{l+1}\left\{(X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(k-1)}})X_{i_{l+1}}(X_{i_{\sigma(k+1)}}\cdots X_{i_{\sigma(l)}})X_{i_{\sigma(k)}}\right.$
$-(X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(k-1)}})[X_{i_{l+1}}~{},~{}(X_{i_{\sigma(k+1)}}\cdots X_{i_{\sigma(l)}})]X_{i_{\sigma(k)}}$
$-\left.(X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(k-1)}})[(X_{i_{\sigma(k+1)}}\cdots X_{i_{\sigma(l)}})X_{i_{l+1}}~{},~{}X_{i_{\sigma(k)}}]~{}\right\}.$ (3.15)
The terms that contain [,][~{},~{}][ , ] can be written as expressions of degree $l$ or lower, and are uniquely symmetrized by the assumption. Therefore, it is sufficient to look at the first term on the right-hand side of (3.15). The first term is written by
$\frac{1}{(l+1)!}\sum_{\sigma\in Sym(l+1)}X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(l+1)}}=X_{(i_{1},\cdots,i_{l+1})}.$
The proposition was thus proved. ∎
To replace $X_{i}$ by $e^{(\mu)}_{i}$ each base $E_{i}$ is expressed as (3.13). Because $E_{i}$’s expression that represents as (3.13) is not uniquely determined, we fix one expression of each $E_{i}$ by some (3.13). Then any $f=\sum_{i}a^{i}E_{i}\in T_{\mu}=\langle e^{(\mu)}\rangle$ is uniquely written by
$f=\sum_{k}^{n_{\mu}}\sum_{i}a^{i}E^{(i)}_{j_{1},\cdots,j_{k}}e^{(\mu)}_{(j_{1},\cdots,j_{k})}=\sum_{k}^{n_{\mu}}f^{(k)}_{(j_{1},\cdots,j_{k})}e^{(\mu)}_{(j_{1},\cdots,j_{k})},$ (3.16)
where $f^{(k)}_{(j_{1},\cdots,j_{k})}=\sum_{i}a^{i}E^{(i)}_{j_{1},\cdots,j_{k}}$. Using this fixed expression of $\{E_{i}\}$, let us introduce a linear map $\phi_{\mu}:T_{\mu}\rightarrow A_{\mathfrak{g}}$ such that
$\phi_{\mu}(f)=\sum_{k}^{n_{\mu}}f^{(k)}_{(j_{1},\cdots,j_{k})}x_{j_{1}}\cdots x_{j_{k}}$ (3.17)
for $f=\sum_{k}^{n_{\mu}}f^{(k)}_{(j_{1},\cdots,j_{k})}e^{(\mu)}_{(j_{1},\cdots,j_{k})}$. In short, for $e^{(\mu)}_{(i_{1},\cdots,i_{k})}$ in (3.13)
$\phi_{\mu}(e^{(\mu)}_{(i_{1},\cdots,i_{k})})=x_{i_{1}}\cdots x_{i_{k}}.$
It is clear from this definition, that
$q_{\mu}\circ\phi_{\mu}=Id$ (3.18)
is satisfied. Using this map $\phi_{\mu}$, we compare how similar the algebra $T_{\mu}$ and the Poisson algebra ${\mathbb{C}}[x]$ as algebras. In this paper, “asymptotic algebra homomorphic map” is defined as follows.
Definition 3.8 (Asymptotic algebra homomorphism).
Let $A$ be a fixed Poisson algebra. Consider a sequence of vector spaces $V^{m}~{}(m=1,2,\cdots)$ and a sequence of subalgebras of $End(V^{m})$ denoted by $B^{m}~{}(m=1,2,\cdots)$. Consider a weak matrix regularization $\{q_{m}:A\to B^{m}\subset End(V^{m})\}$. Each noncommutativity parameter $\hbar(q_{m})$ is abbreviated as $\hbar(m)$. Suppose that there exists a series of linear maps $\{\phi_{m}:B^{m}\to A\}$ and a series of basis of $B^{m}$ satisfying the following conditions.
  1. 1.
    $\{\phi_{m}(X_{i})|\mbox{$X_{i}$ is in the basis of $B^{m}$}\}$ is linearly independent, and each $\phi_{m}(X_{i})$ is independent of $\hbar(m)$.
  2. 2.
    As $\dim V^{m}\rightarrow\infty$, the number of pairs of basis elements $X_{i},X_{j}\in B^{m}$ satisfying the following condition tends infinity.
    $\phi_{m}(X_{i}X_{j})=\phi_{m}(X_{i})\phi_{m}(X_{j})+\tilde{O}(\hbar(m)).$
In this case, we say that $\phi_{m}$ is an asymptotic algebra homomorphism.
The basic idea underlying this definition and the subsequent discussions is presented in [61].
The difficulty in showing that $\phi_{\mu}$, as defined by (3.17), is an asymptotic algebra homomorphism lies in the fact that, in the case of matrix representation, the symmetrized $E_{i}E_{j}$ is not always included in a basis. To observe the property of asymptotic homomorphism of $\phi_{\mu}$, we further restrict the construction method of a basis $E_{i}(i=1,2,\cdots,D)$ as follows. We shall name index multisets as $I^{1}_{j}=\{\!\!\{j\}\!\!\},I^{2}_{j}=\{\!\!\{j_{1},j_{2}\}\!\!\},\cdots,I^{k}_{j}=\{\!\!\{j_{1},\cdots,j_{k}\}\!\!\}$ and denote $e^{(\mu)}_{(j_{1},\cdots,j_{k})}$ by $e^{(\mu)}_{I^{k}_{j}}$. Note that a multiset differs from a set in that it distinguishes the degree of overlap of its elements.
A notation such as $I^{k}_{j}\sqcup I^{l}_{i}:=\{\!\!\{j_{1},\cdots,j_{k},i_{1},\cdots,i_{l}\}\!\!\}$ is also used.
  1. 1.
    We choose the unit matrix $Id$ and $E_{I_{1}^{1}}=e^{(\mu)}_{1},E_{I_{2}^{1}}=e^{(\mu)}_{2},\cdots,E_{I^{1}_{d}}=e^{(\mu)}_{d}$ as a part of the basis. For the sake of uniform description, we denote these as $E_{I^{0}}:=Id~{}(d_{0}:=1)$ and $E_{I^{1}_{i}}(i=1,\cdots,d=:d_{1})$.
  2. 2.
    From $e^{(\mu)}_{(1,1)},e^{(\mu)}_{(1,2)},\cdots,e^{(\mu)}_{(1,d)},e^{(\mu)}_{(2,2)},e^{(\mu)}_{(2,3)},\cdots,e^{(\mu)}_{(2,d)},\cdots e^{(\mu)}_{(d,d)}$ , the maximum linearly independent terms of these are selected and added to $Id=:E_{I^{0}}~{}(d_{0}:=1)$ and $E_{I^{1}_{i}}(i=1,\cdots,d_{1})$ to be used as the basis. Let $d_{2}$ be the number of them, and $E_{I^{2}_{j}}(j=1,\cdots,d_{2})$ be the element of the basis. The index multiset ${I^{2}_{j}}(j=1,\cdots,d_{2})$ is also fixed.
  3. 3.
    We choose the largest number of linearly independent elements from $e^{(\mu)}_{I^{1}_{i}\sqcup I^{2}_{j}}(1\leq i\leq d_{1},1\leq j\leq d_{2})$, and if more linearly independent elements can be made from $e^{(\mu)}_{I^{1}_{i}\sqcup I^{1}_{j}\sqcup I^{1}_{k}}(1\leq i,j,k\leq d_{1})$, they are also added to the basis components. We denote them as $E_{I^{3}_{j}}(j=1,\cdots,d_{3})$. Here $d_{3}$ is the maximum number of linear independent terms. The fixed index multiset ${I^{3}_{j}}(j=1,\cdots,d_{3})$ are elements of $\{I^{1}_{i}\sqcup I^{1}_{j}\sqcup I^{1}_{k}~{}|~{}1\leq i,j,k\leq d\}$.
  4. 4.
    Using $e^{(\mu)}_{I^{2}_{k}\sqcup I^{2}_{j}}(1\leq j,k\leq d_{2})$ and $e^{(\mu)}_{I^{1}_{i}\sqcup I^{3}_{j}}(1\leq i\leq d_{1},1\leq j\leq d_{3})$, we do the similar process. If necessary, $e^{(\mu)}_{I^{1}_{i}\sqcup I^{1}_{j}\sqcup I^{1}_{k}\sqcup I^{1}_{l}}(1\leq i,j,k,l\leq d_{1})$ is also taken into account in the elements of the basis as well. $E_{I^{4}_{j}}(j=1,\cdots,d_{4})$ are chosen.
  5. 5.
    We repeat the same process as above until the basis of $T_{\mu}$ is completed. $d_{0}+d_{1}+\cdots+d_{n_{\mu}}=D$.
The basis of $T_{\mu}$ can be constructed in this way. In other words, it can be constructed so that the index multiset that distinguishes the elements of a basis is the union of the index multisets of the other elements. This is the point of this method of constructing a basis.
The following asymptotic algebra homomorphisms are established in the correspondence of the multisets of their indices.
Proposition 3.9.
Let $E_{I^{k}_{i}}=e^{(\mu)}_{(i_{1},\cdots,i_{k})},E_{I^{l}_{j}}=e^{(\mu)}_{(j_{1},\cdots,j_{l})}$ and $E_{I^{k}_{i}\sqcup I^{l}_{j}}=e^{(\mu)}_{(i_{1},\cdots,i_{k},j_{1},\cdots,j_{l})}$ be elements of basis of $T_{\mu}$. Then the following is obtained.
$\phi_{\mu}(E_{I^{k}_{i}})\phi_{\mu}(E_{I^{l}_{j}})=\phi_{\mu}(E_{I^{k}_{i}}E_{I^{l}_{j}})+\tilde{O}(\hbar({\mu})).$ (3.19)
Proof.
The left-hand side of (3.19) is given as
$\phi_{\mu}(E_{I^{k}_{i}})\phi_{\mu}(E_{I^{l}_{j}})$ $=x_{i_{1}}\cdots x_{i_{k}}x_{j_{1}}\cdots x_{j_{l}}.$
On the other hand, by using (3.3) $E_{I^{k}_{i}}E_{I^{l}_{j}}$ is expressed as
$E_{I^{k}_{i}}E_{I^{l}_{j}}=e^{(\mu)}_{(i_{1},\cdots i_{k},j_{1},\cdots,j_{l})}-\sum_{m<k+l}\sum_{n=1}^{d_{m}}\hbar(\mu)^{k+l-m}c_{I^{m}_{n},(i_{1},\cdots,j_{l})}E_{I^{m}_{n}},$ (3.20)
where $c_{I^{m}_{n},(i_{1},\cdots,j_{l})}$ is a complex number that does not depend on $\hbar(\mu)$. Because $E_{I^{k}_{i}\sqcup I^{l}_{j}}=e^{(\mu)}_{(i_{1},\cdots,i_{k},j_{1},\cdots,j_{l})}$ is contained in the basis, (3.19) leads to the following equation;
$\phi_{\mu}(E_{I^{k}_{i}}E_{I^{l}_{j}})$ $=\phi_{\mu}(e^{(\mu)}_{(i_{1},\cdots i_{k},j_{1},\cdots,j_{l})}-\sum_{m<k+l}\sum_{n=1}^{d_{m}}\hbar(\mu)^{k+l-m}c_{I^{m}_{n},(i_{1},\cdots,j_{l})}E_{I^{m}_{n}})$
$=x_{i_{1}}\cdots x_{i_{k}}x_{j_{1}}\cdots x_{i_{l}}-\sum_{m<k+l}\sum_{n=1}^{d_{m}}\hbar(\mu)^{k+l-m}c_{I^{m}_{n},(i_{1},\cdots,j_{l})}x_{n_{1}}\cdots x_{n_{m}}.$
Since $\sum_{m<k+l}\sum_{n=1}^{d_{m}}$ is a finite sum, the following conclusion is obtained.
$\phi_{\mu}(E_{I^{k}_{i}})\phi_{\mu}(E_{I^{l}_{j}})=\phi_{\mu}(E_{I^{k}_{i}}E_{I^{l}_{j}})+\tilde{O}(\hbar({\mu})).$ (3.21)
More generally, this proposition can be generalized as follows.
Proposition 3.10.
Let $E_{I^{p_{1}}_{i}},E_{I^{p_{2}}_{j}},\cdots,E_{I^{p_{m}}_{k}}$ and $E_{I^{p_{1}}_{i}\sqcup\cdots\sqcup I^{p_{m}}_{k}}$ be elements of basis of $T_{\mu}$. Then the following is obtained
$\phi_{\mu}(E_{I^{p_{1}}_{i}})\cdots\phi_{\mu}(E_{I^{p_{m}}_{k}})=\phi_{\mu}(E_{I^{p_{1}}_{i}}\cdots E_{I^{p_{m}}_{k}})+\tilde{O}(\hbar({\mu})).$ (3.22)
The proof is obtained in the same way as the proof of Proposition 3.9.
Remark.
Consider the case where the dimension ${\rm dim}V^{\mu}$ of the representation space $V^{\mu}$ increases. Suppose that the representation $\rho^{\mu}$ is an irreducible representation. In this case, the dimension of the algebra $T_{\mu}$ generated by the basis of the Lie algebra, i.e., the number of $E_{1},...,E_{D}$, increases. Therefore, the number of pairs of elements of the basis that satisfy the above Proposition 3.9 also increases. In the limit ${\rm dim}V^{\mu}\rightarrow\infty$, we can understand $\mathbb{C}[x]$ as an approximation to the algebra $T_{\mu}$ in the sense that the number of pairs of elements satisfying the above Proposition 3.9 increases infinitely.
Remark.
Using this $\phi_{\mu}$, we can write the necessary condition that a sequence of quantizations $\{q_{\mu}:A_{\mathfrak{g}}\to T_{\mu}\subset End(V^{\mu})\}$ is a weak matrix regularization as
$\phi_{\mu}([q_{\mu}(f),q_{\mu}(g)])=\hbar({\mu})\phi_{\mu}(q_{\mu}(\{f,g\}))+\tilde{O}(\hbar^{2}(\mu))$
for any $f,g\in A_{\mathfrak{g}}$. The existence of this asymptotic algebra homomorphism map $\phi_{\mu}$ clarifies that the quantization correctly reflects noncommutativity. There exists an $\hbar(\mu)$ in the image of $q_{\mu}$ that is derived from the degree of the polynomial in $A_{\mathfrak{g}}$. On the other hand, in the image of this $\phi_{\mu}$, only the $\hbar(\mu)$ derived from commutators or, due to the finite dimensionality of the matrix, the $\hbar(\mu)$ that appears only from higher-degree terms exceeding the matrix dimension (i.e., the $\hbar(\mu)$ caused by Lemma 3.2) appears. The $\hbar(\mu)$ that appears in Lemma 3.2 originates from higher-order terms corresponding to monomials that are mapped outside the adaptive part of $gl(V^{\mu})$. Therefore, such terms are ignored in the commutative limit as well as higher-order of noncommutativity.
Since the quantization by $q_{\mu}$ is defined as $q_{\mu}(x_{i})=e^{(\mu)}_{i}=\hbar(\mu)\rho^{\mu}(e_{i})$, the degree of $\hbar$ changes depending on the degree of the monomial. As already pointed out in the case of the Fuzzy sphere by Chu, Madore, and Steinakker [28], the classical(commutative) limit of matrix regularization changes its result by fixing the ratio of 1/$1/\hbar$ and $\dim V^{\mu}$. The Casimir operator is a quantity that has one typical feature that is consistent with a Lie algebra and a corresponding Lie-Poisson algebra.
Okubo [66] gave the formula of eigenvalue of $k$-th Casimir $C_{k}$;
$C_{k}=\sum_{\mu}\frac{{\dim}V^{\mu}}{{\dim}\mathfrak{g}}\left(\frac{C_{2}(V^{\mu})-2C_{2}(ad)}{2}\right)^{k}.$ (3.23)
Note the omission of the identity matrix and the abuse of symbols for eigenvalues and matrices. This formula corresponds to the eigenvalue of Casimir operator that does not depend on $\hbar$. Casimir polynomials of $A_{\mathfrak{g}}$ is defined by
$\{x_{i},f(x)\}=0~{}(i=1,\cdots,d),$ (3.24)
and we denote the set of all Casimir polynomials of $A_{\mathfrak{g}}$ by $CaP$.
Proposition 3.11.
For any $k$-th degree Casimir polynomial $f^{C}_{k}\in CaP$, $\hbar^{k}(\mu)C_{k}:=q_{\mu}(f^{C}_{k})\in T_{\mu}$ is a Casimir operator.
This has already been given in [1], albeit for enveloping algebras. Also, the proof of the first half of Proposition 4.3 corresponds to that proof, so you may refer to it there. Since (3.23), $\lim_{\dim V^{\mu}\rightarrow\infty}|C_{k}|=\infty$. So, if $\hbar$ is chosen such that the absolute value of the eigenvalue of $q_{\mu}(f^{C}_{k})$
${|C_{k}|}{|\hbar(\mu)|^{k}}$ (3.25)
is fixed, then the correspondent $k$-th Casimir polynomial is survived under the $\hbar(\mu)\rightarrow 0,{\dim}V^{\mu}\rightarrow\infty$. For later convenience, we will uniquely determine $\hbar(\mu)$ as follows. Let $C^{k}_{i}(x)=\sum_{\alpha}C_{i,\alpha}x^{J}=\sum_{\alpha}C_{i,\alpha}x_{j_{1}}\cdots x_{j_{k}}$ be one of the $k$-th Casimir polynomial. We define $C^{k}_{i}(e^{\mu})$ by
$C^{k}_{i}(e^{\mu}):=q_{\mu}(C^{k}_{i}(x))=\sum_{J}C_{i,J}e^{(\mu)}_{(j_{1},\cdots,j_{k})}=\frac{\hbar^{k}(\mu)}{k!}\sum_{J}C_{i,J}\sum_{\sigma\in Sym(k)}\rho^{\mu}(e_{j_{\sigma(1)}})\cdots\rho^{\mu}(e_{j_{\sigma(k)}}).$ (3.26)
From Schur’s lemma, this Casimir operator is proportional to the unit matrix. It can be fixed to any eigenvalue $\lambda^{k}_{i}\in{\mathbb{C}}$ of the matrix $C^{k}_{i}(e^{\mu})$ by determining $\hbar(\mu)$ appropriately;
$C^{k}_{i}(e^{\mu})=\hbar^{k}(\mu)C_{k}=\lambda^{k}_{i},$ (3.27)
for any $q_{\mu}:A_{\mathfrak{g}}\to V^{\mu}$. This is the reason for describing varieties in this paper in terms of Casimir polynomials of degree $k$.
We comment on this eigenvalue (3.27). By “the eigenvalue is fixed”, we mean that it is the same value for any $q_{\mu}$ in the series of matrix regularization $\{q_{\mu}\}$. $\lambda^{k}_{i}$ is still a polynomial of order $k$ in $\hbar(\mu)$ and is still an $\tilde{O}(\hbar^{k}(\mu))$.
4 (Weak) Matrix regularization for Lie-Poisson varieties
Casimir polynomials of $A_{\mathfrak{g}}$ is defined by
$\{x_{i},f(x)\}=0~{}(i=1,\cdots,d),$ (4.1)
and we denote the set of all Casimir polynomials of $A_{\mathfrak{g}}$ by $CaP$. Consider a vector space $C\subset CaP$ whose basis is $\{f_{i}^{C}\}$, i.e., $C:=\{\sum_{i}a_{i}f_{i}^{C}~{}|~{}a_{i}\in\mathbb{C},~{}f_{i}^{C}\in CaP\}$. We introduce an ideal of $A_{\mathfrak{g}}$ generated by $C$ as
$I(C):=\left\{\sum f_{i}^{C}(x)g_{i}(x)\in A_{\mathfrak{g}}~{}|~{}f_{i}^{C}(x)\in C,~{}g_{i}(x)\in A_{\mathfrak{g}}~{}\right\}.$ (4.2)
This ideal is compatible with the Poisson structure because $\{x_{i},f_{j}^{C}(x)\}=0$. So, we can introduce the new Poisson algebra as follows:
$A_{\mathfrak{g}}/I(C):=\left\{[f(x)]~{}\mid~{}f(x)\in A_{\mathfrak{g}}\right\},$ (4.3)
where $[f(x)]=\{f(x)+h(x)~{}\mid h(x)\in I(C)\}$, and the sum and multiplication are defined as $[f(x)]+[g(x)]=[f(x)+g(x)]$ and $[f(x)]\cdot[g(x)]=[f(x)\cdot g(x)]$. The Poisson bracket is also defined by as
$\{[f(x)],[g(x)]\}=[\{f(x),g(x)\}].$
We abbreviate this Poisson algebra $(A_{\mathfrak{g}}/I(C),\cdot,\{~{},~{}\})$ as $A_{\mathfrak{g}}/I(C)$. In this section, we formulate the quantization of this Poisson algebra by means of weak matrix regularization.
Remark.
In this paper, we call $A_{\mathfrak{g}}/I(C)$ “Lie-Poisson variety”. However, it does not mean $I(C)$ is a prime ideal. We use the term variety to mean an algebraic variety defined by $f_{i}^{C}(x)=0$.
The quotient and remainder of a multivariate polynomial cannot be uniquely determined. However, after introducing some monomial ordering like lexicographic ordering and ascertaining the monomial ordering on multivariate polynomials, we can define unique remainder of a polynomial.
It is known that if we fix the ordering every ideal $I$ has a unique reduced Gröbner basis, and the following fact is known. (See the Appendix C for definitions and necessary information on the Gröbner basis.).
Theorem 4.1.
(See for example [33, 30].) Fix a monomial ordering on $K[x_{1},\dots,x_{n}]$.
  1. 1.
    Every ideal $I\subset K[x_{1},\dots,x_{n}]$ has a unique reduced Gröbner basis.
  2. 2.
    Let ${g_{1},\dots,g_{m}}$ be the Gröbner basis for the ideal $I$ in $K[x_{1},\dots,x_{n}]$. Then every polynomial $f\in K[x_{1},\dots,x_{n}]$ can be written uniquely in the form
    $f=h+r$ (4.4)
    where $h\in I$ and no monomial term of the $r$ is divisible by any $LT(g_{i})$.
In the following, we fix a monomial ordering for every Lie-Poisson variety by the graded lexicographic ordering.
4.1 Formulation of quantization via enveloping algebra
In this subsection, we formulate a quantization map from $A_{\mathfrak{g}}/I(C)$ to a quotient algebra of the universal enveloping algebra by a two-sided ideal.
We use $\mathbb{C}[\hbar]$ as the commutative ring $R$, where $\hbar$ is a complex variable in this subsection. As in Subsection 3.2, we use the enveloping algebra $\mathcal{U}_{\mathfrak{g}}[\hbar]$. Recall the two-side ideal is $I=\{\sum_{i,j,k,l}a_{k}(X_{i}X_{j}-X_{j}X_{i}-[X_{i},X_{j}])b_{l}~{}|~{}a_{k},b_{l}\in R\langle X\rangle\}$, and $\mathcal{U}_{\mathfrak{g}}[\hbar]$ is defined by $\mathcal{U}_{\mathfrak{g}}[\hbar]=R\langle X\rangle/I.$
We comment on why we need to introduce $\mathcal{U}_{\mathfrak{g}}[\hbar]$ here. The deformation using the independent parameter $\hbar$ as above means that the generator $X_{i}$ is completely independent of $\hbar$. Fuzzy space or matrix regularization, $q_{\mu}$, is an expansion of $\hbar$, like a Taylor expansion. Care is needed to distinguish the degree of $\hbar$, which represents noncommutativity, and this requires an asymptotic algebraic homomorphism. On the other hand, the quantization developed in this subsection has a major difference in that the degree of the generators and the degree of $\hbar$ are not related.
We are dealing with what has long been known as the canonical mapping from $A_{\mathfrak{g}}$ to $\mathcal{U}_{\mathfrak{g}}$ in the studies by Abellanas-Alonso [1], (see also [32]) , and we replace its target space from $\mathcal{U}_{\mathfrak{g}}$ to $\mathcal{U}_{\mathfrak{g}}[\hbar]$.
Definition 4.2 (Dixmier, Abellanas-Alonso).
We define a canonical linear map $q_{U}:A_{\mathfrak{g}}\to\mathcal{U}_{\mathfrak{g}}[\hbar]$ by
$x_{i_{1}}\cdots x_{i_{k}}\mapsto\frac{1}{k!}\sum_{\sigma\in Sym(k)}X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(k)}}.$ (4.5)
Proposition 4.3.
$q_{U}:A_{\mathfrak{g}}\to\mathcal{U}_{\mathfrak{g}}[\hbar]$ is a quantization i.e.,
$[q_{U}(f),q_{U}(g)]=\hbar~{}q_{U}(\{f,g\})+\tilde{O}(\hbar^{2})$
for $f,g\in A_{\mathfrak{g}}$. In other words, $q_{U}\in Q$. Especially, if $\min\{\deg f,\deg g\}\leq 1$, then
$[q_{U}(f),q_{U}(g)]=\hbar~{}q_{U}(\{f,g\}).$ (4.6)
Proof.
Since there is an anti-symmetry with respect to the permutation of $f$ and $g$, it is enough that we write the proof for $\deg f\leq\deg g$. In addition, since it is sufficient to prove the case of monomials by using the linearity of $q_{U}$, Poisson brackets, and Lie brackets, we treat $f$ and $g$ as monomials below. The case of $\deg f=0$ is trivial. Let us consider the case of $\deg f=1$. For monomials $f(x)=x_{i}$ and $g=x^{I^{m}}=x_{i_{1}}\cdots x_{i_{m}}$ ,
$q_{U}(\{x_{i},x^{I^{m}}\})$ $=\sum_{l,k}f^{k}_{il}q_{U}(x_{k}\partial_{l}x^{I^{m}})$
$=\sum_{l,k}f^{k}_{il}~{}q_{U}\left(x_{k}\sum_{n=1}^{m}x_{i_{1}}\cdots x_{i_{n-1}}\delta_{\beta_{n}l}x_{i_{n+1}}\cdots x_{i_{m}}\right).$
Let $J^{(n)}$ be the indices $(j_{1}^{(n)},j_{2}^{(n)},\cdots,j_{m}^{(n)}):=(i_{1},\cdots,i_{n-1},k,i_{n+1},\cdots,i_{m})$. Then, we obtain
$q_{U}(\{x_{i},x^{I^{m}}\})=\frac{1}{m!}\sum_{\sigma\in Sym(m)}\sum_{n=1}^{m}\sum_{k=1}^{d}\left(X_{j_{\sigma(1)}^{(n)}}\cdots(f^{k}_{ii_{n}}X_{j_{\sigma(\sigma^{-1}(n))}^{(n)}})\cdots X_{j_{\sigma(m)}^{(n)}}\right).$
Here $f^{k}_{ii_{n}}X_{j_{\sigma(\sigma^{-1}(n))}^{(n)}}\!\!\!\!\!\!\!=f^{k}_{ii_{n}}X_{j_{n}^{(n)}}=f^{k}_{ii_{n}}X_{k}$ is the $\sigma^{-1}(n)$-th item from the left. Finally, we find
$\hbar~{}q_{U}(\{x_{i},x^{I^{m}}\})$ $=\frac{1}{m!}\sum_{\sigma\in Sym(m)}\sum_{n=1}^{m}X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(\sigma^{-1}(n)-1)}}[X_{i},X_{i_{n}}]X_{i_{\sigma(\sigma^{-1}(n)+1)}}\cdots X_{i_{\sigma(m)}}$
$=\frac{1}{m!}\sum_{\sigma\in Sym(m)}\sum_{n=1}^{m}X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(n-1)}}[X_{i},X_{i_{\sigma(n)}}]X_{i_{\sigma(n+1)}}\cdots X_{i_{\sigma(m)}}$
$=\frac{1}{m!}\sum_{\sigma\in Sym(m)}[X_{i},X_{i_{\sigma(1)}}\cdots X_{i_{\sigma(m)}}]=[q_{U}(x_{i}),q_{U}(x^{I^{m}})].$
Therefore, (4.6) is shown for $\min\{\deg f,\deg g\}\leq 1$.
Next, we consider the case with $\deg f\geq 2$. Essentially, the case $\deg f\geq 3$ is the same as the case $\deg f=2$, so for simplicity we only describe the case where $\deg f=2$. For monomials $f(x)=x_{i}x_{j}$ and $g=x^{I^{m}}=x_{i_{1}}\cdots x_{i_{m}}$,
$q_{U}(\{x_{i}x_{j},g\})=q_{U}(x_{i}\{x_{j},g\}+x_{j}\{x_{i},g\})=\sum_{l,k}\left\{q_{U}(x_{i}f^{k}_{jl}x_{k}\partial_{l}x^{I^{m}})+q_{U}(x_{j}f^{k}_{il}x_{k}\partial_{l}x^{I^{m}})\right\}.$ (4.7)
Let us take a closer look at the first term $q_{U}(x_{i}f^{k}_{jl}x_{k}\partial_{l}x^{I^{m}})$. After the similar calculations as in the case of $\deg f=1$, we use $\alpha^{(n,m+1)}$ be the indices $(\alpha_{1},\alpha_{2},\cdots,\alpha_{m+1}):=(i_{1},\cdots,i_{n-1},k,i_{n+1},\cdots,i_{m},i)$.
$\sum_{l,k}\hbar q_{U}(x_{i}f^{k}_{jl}x_{k}\partial_{l}x^{I^{m}})=$
$\frac{1}{(m+1)!}\sum_{\sigma\in Sym(m+1)}\sum_{n=1}^{m}X_{\alpha_{\sigma(1)}}\cdots X_{\alpha_{\sigma(\sigma^{-1}(n)-1)}}[X_{j},X_{i_{n}}]X_{\alpha_{\sigma(\sigma^{-1}(n)+1)}}\cdots X_{i}\cdots X_{\alpha_{\sigma(m+1)}}.$
Here, $X_{i}$ is located at the $\sigma^{-1}(m+1)$-th position from the left. Using the commutation relation, we can write the above with $X_{i}$ at the beginning and at the end as follows.
$\sum_{l,k}\hbar q_{U}(x_{i}f^{k}_{jl}x_{k}\partial_{l}x^{I^{m}})=$ $\frac{1}{2}\frac{m+1}{(m+1)!}X_{i}\!\!\!\!\sum_{\sigma\in Sym(m)}\!\!\sum_{n}X_{i_{\sigma(1)}}\cdots[X_{j},X_{i_{n}}]\cdots X_{i_{\sigma(m)}}$ (4.8)
$+$ $\frac{1}{2}\frac{m+1}{(m+1)!}\!\!\!\!\sum_{\sigma\in Sym(m)}\!\!\sum_{n}X_{i_{\sigma(1)}}\cdots[X_{j},X_{i_{n}}]\cdots X_{i_{\sigma(m)}}X_{i}+\tilde{O}(\hbar^{2}).$
The second term in (4.7) is obtained by swapping the $i$ and $j$ in (4.8). Therefore,
$\hbar~{}q_{U}(\{x_{i}x_{j},g\})=$ $\frac{1}{2}\left(X_{i}[X_{j},q_{U}(x^{I^{m}})]+[X_{j},q_{U}(x^{I^{m}})]X_{i}+X_{j}[X_{i},q_{U}(x^{I^{m}})]+[X_{i},q_{U}(x^{I^{m}})]X_{j}\right)$
$+\tilde{O}(\hbar^{2}).$ (4.9)
On the other hand,
$[q_{U}(x_{i}x_{j}),q_{U}(x^{I^{m}})]=\frac{1}{2}[X_{i}X_{j}+X_{j}X_{i},q_{U}(x^{I^{m}})]$ (4.10)
$=\frac{1}{2}\left(X_{i}[X_{j},q_{U}(x^{I^{m}})]+[X_{j},q_{U}(x^{I^{m}})]X_{i}+X_{j}[X_{i},q_{U}(x^{I^{m}})]+[X_{i},q_{U}(x^{I^{m}})]X_{j}\right).$
From (4.9) and (4.10), the desired result is proven. ∎
Next, in order to construct the quantization of $A_{\mathfrak{g}}/I(C)$, the following ideal $I(C(X))\subset\mathcal{U}_{\mathfrak{g}}[\hbar]$ is also introduced into the enveloping algebra.
$I(C(X))$ :=$:=\left\{\sum a(X)f_{i}^{C}(X)b(X)\in\mathcal{U}_{\mathfrak{g}}[\hbar]~{}|~{}f_{i}^{C}(x)\in CaP,~{}a(X),b(X)\in A_{\mathfrak{g}}~{}\right\}.$ (4.11)
Here, we use
$f_{i}^{C}(X):=q_{U}(f_{i}^{C}(x)),\quad\quad f_{i}^{C}(x)\in CaP:=\{f(x)\in A_{\mathfrak{g}}~{}|~{}\{x_{i},f(x)\}=0(i=1,\cdots,d)\}.$ (4.12)
Note that from (4.6),
$[X_{j},f_{i}^{C}(X)]=[X_{j},q_{U}(f_{i}^{C}(x))]=0,\quad(i=j,\cdots,d).$ (4.13)
In short $f_{i}^{C}(X)$ is a Casimir operator in $\mathcal{U}_{\mathfrak{g}}[\hbar]$. Therefore, it is possible to define $\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$ while keeping it compatible with the commutation relations. We denote $\{f(X)+h(X)~{}|~{}f(X)\in\mathcal{U}_{\mathfrak{g}}[\hbar],h(X)\in I(C(X))\}$ by $[f(X)]$. The sum and product of the algebra $\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$ are defined by $[f]+[g]:=[f+g]$ and $[f][g]=[fg]$, and the commutator product is determined by $[[f],[g]]=[[f,g]]$. Let us generalize $q_{U}:A_{\mathfrak{g}}\to\mathcal{U}_{\mathfrak{g}}[\hbar]$ to $q_{U/I}:A_{\mathfrak{g}}/I(C)\to\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$.
Definition 4.4.
A linear map
$q_{U/I}:A_{\mathfrak{g}}/I(C)\to\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$
is defined as follows. Let $G$ be a reduced Gröbner basis of $I(C)$. For any $f(x)\in A_{\mathfrak{g}}$, $r_{f,G}$ is uniquely determined by Theorem 4.1 as
$f(x)=r_{f,G}(x)+h_{f},\quad h_{f}\in I(C).$
For $\forall[f(x)]\in A_{\mathfrak{g}}/I(C)$, we define
$q_{U/I}([f(x)]):=[q_{U}(r_{f,G}(x))],$
where $q_{U}(r_{f,G}(x))$ is determined by Definition 4.2.
Theorem 4.5.
The above $q_{U/I}:A_{\mathfrak{g}}/I(C)\to\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$ is a quantization;
$[q_{U/I}([f]),q_{U/I}([g])]=\hbar q_{U/I}(\{[f],[g]\})+\tilde{O}(\hbar^{2}),$ (4.14)
for $\forall[f],[g]\in A_{\mathfrak{g}}/I(C)$. $\tilde{O}(\hbar^{n})$ is used in the sense of Example A.4 in Appendix A.
Proof.
In the following, for any polynomial $f$ and subscript $i$, we write $f:=h_{i}+r_{i}$ to mean equation (4.4), i.e., $h_{i}\in I(C),r_{i}\notin I(C)$. For any $f,g\in A_{\mathfrak{g}}$, from
$\{f,g\}=h_{\{f,g\}}+r_{\{f,g\}},$
and
$\{f,g\}=\{r_{f}+h_{f},r_{g}+h_{g}\}=\{r_{f},r_{g}\}+h,$
where $f=h_{f}+r_{f},~{}g=h_{g}+r_{g}$ and $h\in I(C)$, we obtain
$\{r_{f},r_{g}\}=r_{\{f,g\}}+h_{rfg},\quad r_{\{r_{f},r_{g}\}}=r_{\{f,g\}}.$ (4.15)
Here the uniqueness described in Theorem 4.1 is used to obtain the above result. Therefore we find
$\hbar\,q_{U/I}([\{f,g\}])=\hbar\,[q_{U}(r_{\{f,g\}})]=\hbar\left[q_{U}(\{r_{f},r_{g}\})-q_{U}(h_{rfg})\right].$ (4.16)
From the Proposition 4.3,
$\left[\hbar\,q_{U}(\{r_{f},r_{g}\})\right]$ $=\left[[q_{U}(r_{f}),q_{U}(r_{g})]\right]+\tilde{O}(\hbar^{2})$
$=[q_{U/I}([r_{f}]),q_{U/I}([r_{g}])]+\tilde{O}(\hbar^{2})=[q_{U/I}([f]),q_{U/I}([g])]+\tilde{O}(\hbar^{2})$ (4.17)
From (4.16) and (4.17), this proof is complete if we can show that
$[q_{U}(h_{rfg})]=\tilde{O}(\hbar).$
Recall that $h_{rfg}$ is an element of $I(C)$ that is generated by Casimir polynomials;
$h_{rfg}=\sum f_{i}^{C}(x)k_{i}(x),$
where $k_{i}(x)\in A_{\mathfrak{g}}$ and $f_{i}^{C}(x)$ is a Casimir polynomial. So, it is enough that we show
$[q_{U}(C(x)x^{I})]=\tilde{O}(\hbar),$
for any Casimir polynomial $C(x)$ and any monomial $x^{I}=x_{i_{1}}\cdots x_{i_{m}}$. Recall that $q_{U}(C(x))$ is also a Casimir operator as we saw in (4.13). So, $q_{U}(C(x))q_{U}(x^{I})$ is in the ideal $IC(X)$, and
$[q_{U}(C(x)x^{I})]=[q_{U}(C(x)x^{I})-q_{U}(C(x))q_{U}(x^{I})].$ (4.18)
We put $C(x)=\sum_{J}C_{J}x^{J}~{}(C^{k}_{J}\in{\mathbb{C}},x^{J}=x_{j_{1}}\cdots x_{j_{k}})$, then
$q_{U}(C(x)x^{I})=\sum_{J}C_{J}q_{U}(x^{J}x^{I}).$
By the similar discussions with the proof for Lemma 3.7,
$q_{U}(x^{J}x^{I})=X_{(j_{i},\cdots,j_{k},i_{1},\cdots,i_{m})}=X_{(j_{i},\cdots,j_{k})}X_{(i_{1},\cdots,i_{m})}+\tilde{O}(\hbar)=q_{U}(x^{J})q_{U}(x^{I})+\tilde{O}(\hbar).$
Using this, finally (4.18) is written as
$[q_{U/I}(C(x)x^{I})]$ $=\left[\sum_{J}C_{J}q_{U}(x^{J}x^{I})-q_{U}(C(x))q_{U}(x^{I})\right]$
$=\left[\sum_{J}C_{J}(q_{U}(x^{J}x^{I})-q_{U}(x^{J})q_{U}(x^{I}))\right]$
$=[\tilde{O}(\hbar)]=\tilde{O}(\hbar).$ (4.19)
Here $[\tilde{O}(\hbar)]=\tilde{O}(\hbar)$ is obtained from Definition A.2.
Remark.
As you can see from the above proof, there is no quantization $A_{\mathfrak{g}}/I(C)\to\mathcal{U}_{\mathfrak{g}}[\hbar]$ in this construction method. Therefore, it is necessary to discuss the quantization $q_{U/I}:A_{\mathfrak{g}}/I(C)\to\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X)).$
4.2 (Weak) Matrix regularization for Lie-Poisson varieties $A_{\mathfrak{g}}/I(C)$
Up to the previous subsection, we have discussed Casimir polynomials without any conditions. In this subsection, we will impose restrictions so that it becomes Casimir relations that we need. This will make it possible to construct matrix regularization of Lie-Poisson varieties. As already mentioned around (3.25), by balancing the rate at which $\hbar$ approaches zero with the rate at which the dimension of the representation matrix diverges, one can ensure that Casimir polynomials of degree $k$ define a Lie-Poisson variety in the classical(commutative) limit. In this following, $k<n_{\mu}$ is assumed.
We fix $\hbar$ of $\mathcal{U}_{\mathfrak{g}}[\hbar]$ as $\hbar=\hbar(\mu)$. From $\mathcal{U}_{\mathfrak{g}}[\hbar]$ to $gl(V^{\mu})$ there is a representation $\rho_{U\mu}$ as an algebra homomorphism defined by
$\rho_{U\mu}(X_{i}):=e^{(\mu)}_{i}.$ (4.20)
When we fix $\hbar=\hbar(\mu)$, using this representation we obtain $k$-th degree Casimir operator $C^{k}_{i}(e^{\mu}):=\rho_{U\mu}\circ q_{U}(C^{k}_{i}(x))=q_{\mu}(C^{k}_{i}(x))$. Here $C^{k}_{i}(x)$ is a Casimir polynomial of degree $k$. It is clear from the definition of $\rho_{U\mu}$ that the commutation relation is unchanged even for the image of $\rho_{U\mu}$, so $C^{k}_{i}(e^{\mu})$ is a Casimir operator. If it is an irreducible representation, then $C^{k}_{i}(e^{\mu})$ is proportional to the unit matrix if it is not zero, since Schur’s lemma. In the irreducible representation, the Casimir operator can be characterized by an eigenvalue. Let $\lambda^{k}_{i}$ denote the eigenvalues of the Casimir operators of $k$-th degree $C^{k}_{i}(e^{\mu})$. Let us consider the representation for constructing the matrix regularization in that case. So we chose the generators of the ideal $I(C)\subset A_{\mathfrak{g}}$ as
$(f_{i}^{C}(x):=C^{k}_{i}(x)-\lambda^{k}_{i})_{i\in\{1,\cdots,l\}},$ (4.21)
where $l\in{\mathbb{N}}$ is chosen as a number of equations to determine a Lie-Poisson variety. The possible values of $l$ are also restricted by the Lie algebra $\mathfrak{g}$. (For example, in the case of $\mathfrak{su}(n)$, $l$ is at most $1$.) Let us reconstruct all in Subsection 4.1 using
$I(C):=\left\{\sum f_{i}^{C}(x)g_{i}(x)\in A_{\mathfrak{g}}~{}|~{}f_{i}^{C}(x)=C^{k}_{i}(x)-\lambda^{k}_{i},~{}g_{i}(x)\in A_{\mathfrak{g}}~{}\right\}\subset A_{\mathfrak{g}}.$ (4.22)
Then,
$I(C(X))$ :=$:=\left\{\sum a_{i}(X)f_{i}^{C}(X)b_{i}(X)\in\mathcal{U}_{\mathfrak{g}}[\hbar]~{}|~{}a_{i}(X),b_{i}(X)\in\mathcal{U}_{\mathfrak{g}}[\hbar]~{}\right\}.$ (4.23)
Here, $f_{i}^{C}(X)$ is a Casimir operator in $\mathcal{U}_{\mathfrak{g}}[\hbar]$ given as
$f_{i}^{C}(X):=q_{U}(f_{i}^{C}(x))=q_{U}(C^{k}_{i}(x)-\lambda^{k}_{i}).$ (4.24)
$\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X)),q_{U/I}$ and so on are defined by using these ideals.
Definition 4.6.
A linear function $\rho_{U/I,\mu}:\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))\to gl(V^{\mu})$ is defined as follows. For any monomial $X^{I}=X_{i_{1}}\cdots X_{i_{m}}~{}(m\in{\mathbb{N}})$ in $\mathcal{U}_{\mathfrak{g}}[\hbar]$,
$\rho_{U/I,\mu}([X^{I}])=e^{(\mu)}_{i_{1}}\cdots e^{(\mu)}_{i_{m}}.$ (4.25)
Let us check the consistency of this definition of $\rho_{U/I,\mu}$. Let $\Sigma$ be the set of generators of all relations,
$\Sigma:=\left\{X_{j}X_{k}-X_{k}X_{j}-[X_{j},X_{k}],~{}q_{U}(C^{k}_{i}(x))-\lambda^{k}_{i}~{}|~{}1\leq j,k\leq d,i=1,2,\cdots,l\right\}.$
We denote the ideal generated by $\Sigma$ by
:=${\mathfrak{I}}:=\left\{\sum_{i,j,k}a_{i}{\mathcal{I}}_{j}b_{k}~{}|~{}a_{i},b_{k}\in R\langle X\rangle,~{}{\mathcal{I}}_{j}\in\Sigma\right\}.$
In other words, $\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))=R\langle X\rangle/{\mathfrak{I}}$. For any $[f(X)]=[g(X)]$, there exists $h(X)\in{\mathfrak{I}}$ such that $g(X)=f(X)+h(X)$. Note that
$\rho_{U/I,\mu}(X_{j}X_{k}-X_{k}X_{j}-[X_{j},X_{k}])$ $=e^{(\mu)}_{j}e^{(\mu)}_{k}-e^{(\mu)}_{k}e^{(\mu)}_{j}-[e^{(\mu)}_{j},e^{(\mu)}_{k}]=0,$ (4.26)
$\rho_{U/I,\mu}(q_{U}(C^{k}_{i}(x))-\lambda^{k}_{i})$ $=0.$ (4.27)
These equations mean $h(e^{(\mu)})=0$, so we obtain
$\rho_{U/I,\mu}([g(X)])=g(e^{(\mu)})=f(e^{(\mu)})+h(e^{(\mu)})=f(e^{(\mu)})=\rho_{U/I,\mu}([f(X)]).$ (4.28)
In this way, it was confirmed that $\rho_{U/I,\mu}:\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))\to gl(V^{\mu})$ is well-defined. In addition, $\rho_{U/I,\mu}$ is apparently algebra homomorphism. Linearity is trivial from the definition, and the product is as follows.
$\rho_{U/I,\mu}([f(X)][g(X)])=\rho_{U/I,\mu}([f(X)g(X)])=f(e^{(\mu)})g(e^{(\mu)})=\rho_{U/I,\mu}([f(X)])\rho_{U/I,\mu}([g(X)]).$
Lemma 4.7.
For the algebra homomorphism $\rho_{U/I,\mu}:\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))\to gl(V^{\mu})$, if $[f(X)]\in\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$ is $\tilde{O}(\hbar^{n})$ in the sense of Example A.4 in Appendix A, then $\rho_{U/I,\mu}([f(X)])$ is $\tilde{O}(\hbar^{n})$.
Proof.
When $[f(X)]=\tilde{O}(\hbar^{n})$, there exists $h(X)\in\mathcal{U}_{\mathfrak{g}}[\hbar]$ such that $f(X)=h(X)+I(X)~{}(I(X)\in{\mathfrak{I}})$ and $h(X)=\tilde{O}(\hbar^{n})$ by its definition. $h(X)=\tilde{O}(\hbar^{n})$ means that every $a_{J}(\hbar)\in{\mathbb{C}}[\hbar]$ in $h(X)=\sum_{J}a_{J}(\hbar)X^{J}$ satisfies
$\lim_{x\rightarrow 0}\left|\frac{a_{J}(x\hbar)}{x^{n}}\right|<\infty.$
Recall that $I(e^{(\mu)})=0$, then
$\rho_{U/I,\mu}([f(X)])$ $=h(e^{(\mu)})=\sum_{J}a_{J}(\hbar)e^{(\mu)}_{j_{1}}\cdots e^{(\mu)}_{j_{m}}$
$=\sum_{J}\hbar^{|J|}a_{J}(\hbar)\rho^{\mu}(e_{j_{1}})\cdots\rho^{\mu}(e_{j_{m}}).$ (4.29)
Here we denote $|J|:=m$ for $J=(j_{1},j_{2},\cdots,j_{m})$. Then, we find
$\rho_{U/I,\mu}([f(X)])$ $=\sum_{J}\tilde{O}(\hbar^{n+|J|})=\tilde{O}(\hbar^{n}).$
Recall that $\hbar$ in $\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$ was a variable introduced independently of $\hbar(\mu)$, at first. Using the discussions in Subsection 4.1 and this $\rho_{U/I,\mu}$ with Lemma 4.7, we get the following theorem.
Theorem 4.8.
The linear map $q^{pre}_{\mu}:A_{\mathfrak{g}}/I(C)\to gl(V^{\mu})$ defined by
$q^{pre}_{\mu}:=\rho_{U/I,\mu}\circ q_{U/I}$ (4.30)
is a weak matrix regularization, i.e., $\forall[f],[g]\in A_{\mathfrak{g}}/I(C)$ there exists $P_{\mu}=\sum_{i}^{D}c_{i}(\hbar(\mu))E_{i}\in T_{\mu}$, where $c_{i}(\hbar)$ is a polynomial in $\hbar(\mu)$, such that
$[q^{pre}_{\mu}([f]),q^{pre}_{\mu}([g])]=\hbar q^{pre}_{\mu}(\{[f],[g]\})+\hbar^{2}(\mu)P_{\mu}.$ (4.31)
The $q^{pre}_{\mu}:A_{\mathfrak{g}}/I(C)\to gl(V^{\mu})$ introduced above can be said to be sufficiently weak matrix regularization. Here, we consider imposing the restriction that the degree of the polynomial in the domain of definition, excluding the kernel of the matrix regularization, is less than or equal to $n_{\mu}$. This restriction makes $q^{pre}_{\mu}$ be the generalization of the matrix regularization in the case of $q_{\mu}:A_{\mathfrak{g}}\to gl(V^{\mu})$ or fuzzy sphere.
We introduce a projection map $R_{\mu}:gl(V^{\mu})[\hbar(\mu)]\to gl(V^{\mu})[\hbar(\mu)]$ as follows. Here $\hbar(\mu)$ is considered to be a variable. (Note that $\hbar(\mu)$ was chosen to satisfy $q_{U}(C^{k}_{i}(x))=\lambda^{k}_{i}Id$. The eigenvalue of the Casimir operator, $\lambda^{k}_{i}$, can also be freely chosen by the scaling of $\hbar(\mu)$. In this sense, we use $\hbar(\mu)$ as a variable.) Any $M(\hbar(\mu))\in gl(V^{\mu})[\hbar(\mu)]$ is expressed as $M(\hbar(\mu))=\sum_{0\leq k}\hbar(\mu)^{k}M_{k}=\sum_{0\leq k\leq n_{\mu}}\hbar(\mu)^{k}M_{k}+\tilde{O}(\hbar(\mu)^{n_{\mu}+1})$, where each $M_{k}\in gl(V^{\mu})$ does not depend on $\hbar(\mu)$. For any $M(\hbar(\mu))$, we define $R_{\mu}$ by
$R_{\mu}(M):=\sum_{0\leq k\leq n_{\mu}}\hbar(\mu)^{k}M_{k}.$ (4.32)
Using this $R_{\mu}$, let us define a matrix regularization for $A_{\mathfrak{g}}/I(C)$.
Definition 4.9.
We call
$q_{A/I,\mu}:=R_{\mu}\circ q^{pre}_{\mu}=R_{\mu}\circ\rho_{U/I,\mu}\circ q_{U/I}:A_{\mathfrak{g}}/I(C)\to gl(V^{\mu})$ (4.33)
a quantization of $A_{\mathfrak{g}}/I(C)$.
Theorem 4.10.
$q_{A/I,\mu}$ is a weak matrix regularization, i.e., for $\forall f,g\in A_{\mathfrak{g}}$, there exists $P=\sum_{i}^{D}c_{i}(\hbar(\mu))E_{i}\in T_{\mu}$, where each $c_{i}(\hbar(\mu))$ is a polynomial in $\hbar(\mu)$, such that
$[q_{A/I,\mu}([f]),q_{A/I,\mu}([g])]=\hbar(\mu)q_{A/I,\mu}(\{[f],[g]\})+\hbar^{2}(\mu)P.$ (4.34)
Proof.
Since it is linear, it is sufficient to show the case of monomials. Recall that we fix a monomial ordering for every Lie-Poisson variety by the graded lexicographic ordering. $\forall f,g\in A_{\mathfrak{g}}$ with a reduced Gröbner basis of $I(C)$ is uniquely expressed as $f=r_{f}+h_{f},g=r_{g}+h_{g}$, where $h_{i}\in I(C)$ and $r_{i}\notin I(C)$ for any $i\in A_{\mathfrak{g}}$. We split $q^{pre}_{\mu}([r_{i}])$ as
$q^{pre}_{\mu}([r_{f}])$ $=F_{1}+F_{2},$
$q^{pre}_{\mu}([r_{g}])$ $=G_{1}+G_{2},$
$[F_{1},G_{1}]$ $=\hbar(\mu)(FG_{1}+FG_{2})$
where $\deg F_{1}\leq n_{\mu},\deg G_{1}\leq n_{\mu},\deg FG_{1}\leq n_{\mu}$, and the degree of any monomial in $F_{2}$, $G_{2}$ and $FG_{2}$ is greater than $n_{\mu}$. Using this notation,
$[q_{A/I,\mu}([f]),q_{A/I,\mu}([g])]=[F_{1},G_{1}]=\hbar(\mu)(FG_{1}+FG_{2}).$ (4.35)
On the other hand,
$\hbar(\mu)q_{A/I,\mu}(\{[f],[g]\})$ $=\hbar(\mu)R_{\mu}\circ q^{pre}_{\mu}(\{[r_{f}],[r_{g}]\})$
$=\hbar(\mu)R_{\mu}(\frac{1}{\hbar(\mu)}[q^{pre}_{\mu}([r_{f}]),q^{pre}_{\mu}([r_{g}])])+\hbar^{2}(\mu)P_{1},$ (4.36)
where we use Theorem 4.8. We denote elements in $T_{\mu}$ of the same type as $P$ by $P_{i}=\sum_{j}^{D}c_{j}^{i}(\hbar(\mu))E_{j}$ for $i=1,2$. ($P_{2}$ will be used soon.) There is a notation of $1/\hbar(\mu)$, but this will not cause any misunderstanding because it cancels out with the $\hbar(\mu)$ that arises from the commutator. The first term in the righ-hand side of (4.36) is written as
$R_{\mu}(\frac{1}{\hbar(\mu)}[q^{pre}_{\mu}([r_{f}]),q^{pre}_{\mu}([r_{g}])])$ $=R_{\mu}\left(\frac{1}{\hbar(\mu)}([F_{1},G_{1}]+[F_{1},G_{2}]+[F_{2},G_{1}]+[F_{2},G_{2}])\right)$ (4.37)
If any of $[F_{1},G_{2}],[F_{2},G_{1}]$, or $[F_{2},G_{2}]$ is not 00, then its degree is greater than or equal to $n_{\mu}+2$ and thus $R_{\mu}\big{(}\frac{1}{\hbar(\mu)}([F_{1},G_{2}]+[F_{2},G_{1}]+[F_{2},G_{2}])\big{)}=0$. Here, we used the fact that the commutator product does not change the degree of $\hbar(\mu)$. Therefore,
$R_{\mu}(\frac{1}{\hbar(\mu)}[q^{pre}_{\mu}([r_{f}]),q^{pre}_{\mu}([r_{g}])])$ $=R_{\mu}(\frac{1}{\hbar(\mu)}[F_{1},G_{1}])=FG_{1}.$ (4.38)
From (4.35), (4.36), and (4.38),
$[q_{A/I,\mu}([f]),q_{A/I,\mu}([g])]-\hbar(\mu)q_{A/I,\mu}(\{[f],[g]\})=\hbar(\mu)FG_{2}-\hbar^{2}(\mu)P_{1}.$ (4.39)
Here $FG_{2}=\tilde{O}(\hbar^{n_{\mu}+1}(\mu))$. From Lemma 3.2, $FG_{2}$ is expressed as $\hbar(\mu)P_{2}$. Then we find that (4.34) is satisfied. ∎
For $f=h+r_{f,G}$ with $r_{f,G}=\sum_{J}a_{J}x^{J}=\sum_{\deg x^{J}\leq n_{\mu}}\!\!\!\!\!a_{J}x^{J}+\sum_{\deg x^{J}>n_{\mu}}\!\!\!\!\!a_{J}x^{J}$, the explicit calculation of $q_{A/I,\mu}:A_{\mathfrak{g}}/I(C)\to gl(V^{\mu})$ is given as
$q_{A/I,\mu}([f(x)])=\sum_{m\leq n_{\mu}}\!a_{J}~{}\rho_{U/I,\mu}([X_{(j_{1},\cdots,j_{m})}])=\sum_{m\leq n_{\mu}}\!a_{J}~{}e_{(j_{1},\cdots,j_{m})}^{(\mu)}.$ (4.40)
As can be easily seen from the definition of $q_{A/I,\mu}$, when we chose $I(C)=\{0\}$, it is the same as $q_{\mu}$. Therefore, $q_{A/I,\mu}$ is a generalization of $q_{\mu}$ in Seection 3.2.
At the end of this section, let us show that, as in Proposition 3.9, there is an asymptotic homomorphism between algebra $A_{\mathfrak{g}}/I(C)$ and algebra $T_{\mu}$ with a fixed basis $E_{1},\cdots,E_{D}$. We define a linear map $[\phi_{\mu}]:T_{\mu}\to A_{\mathfrak{g}}/I(C)$ by
$[\phi_{\mu}](e^{(\mu)}_{(i_{1},\cdots,i_{k})}):=[x_{i_{1}}\cdots x_{i_{k}}]$ (4.41)
for $E_{I}=e^{(\mu)}_{(i_{1},\cdots,i_{k})}$.
Then the following is obtained.
Proposition 4.11.
Let $E_{I^{k}_{i}}=e^{(\mu)}_{(i_{1},\cdots,i_{k})},E_{I^{l}_{j}}=e^{(\mu)}_{(j_{1},\cdots,j_{l})}$ and $E_{I^{k}_{i}\sqcup I^{l}_{j}}=e^{(\mu)}_{(i_{1},\cdots,i_{k},j_{1},\cdots,j_{l})}$ be elements of basis of $T_{\mu}$. Then,
$[\phi_{\mu}](E_{I^{k}_{i}})[\phi_{\mu}](E_{I^{l}_{j}})=[\phi_{\mu}](E_{I^{k}_{i}}E_{I^{l}_{j}})+\tilde{O}(\hbar({\mu})).$ (4.42)
Since $[x^{\alpha}][x^{\beta}]=[x^{\alpha}x^{\beta}]$, the proof is the same for Proposition 3.9. Proposition 3.10 can also be generalized to the current case, and is obtained by replacing $\phi_{\mu}$ with $[\phi_{\mu}]$.
Proposition 4.12.
Let $E_{I^{p_{1}}_{i}},E_{I^{p_{2}}_{j}},\cdots,E_{I^{p_{m}}_{k}}$ and $E_{I^{p_{1}}_{i}\sqcup\cdots\sqcup I^{p_{m}}_{k}}$ be elements of basis of $T_{\mu}$ Then the following is obtained
$[\phi_{\mu}](E_{I^{p_{1}}_{i}})\cdots[\phi_{\mu}](E_{I^{p_{m}}_{k}})=[\phi_{\mu}](E_{I^{p_{1}}_{i}}\cdots E_{I^{p_{m}}_{k}})+\tilde{O}(\hbar({\mu})).$ (4.43)
5 Examples
Let us see the examples of weak matrix regularization constructed in Section 4. As the Lie algebra, we consider a semisimple Lie algebra. It corresponds to a classical solution of the mass-deformed IKKT matrix model. We suppose a Lie-Poisson algebra as a classical space when the Lie algebra is regarded as a quantized space (fuzzy space). $\mathfrak{su}(n)$ is a typical example of semisimple Lie algebra. In this section, $\mathfrak{su}(2)$ and $\mathfrak{su}(3)$ are examined. The $\mathfrak{su}(2)$ case is a well-known example of the fuzzy sphere, whereas the other cases provide examples of matrix regularizations that have not been previously studied.
5.1 $\mathfrak{su}(2)$ ; Fuzzy ${\mathbb{R}}^{3}$ and fuzzy sphere
The fuzzy sphere is considered in [41, 61]. See [41, 61, 10, 21] for details. In [71], more general and mathematically precise statements are given. In this subsection, we reconstruct the fuzzy sphere using the method for constructing the weak matrix regularization of a Lie-Poisson variety in Section 4 of this paper. In other words, we confirm that the matrix regularization in this paper is a generalization of the method for constructing the fuzzy sphere to the case of an arbitrary Lie-Poisson variety.
Let us consider $\mathfrak{su}(2)$ as $\mathfrak{g}$. The enveloping algebra of $\mathfrak{su}(2)$, $\mathcal{U}_{\mathfrak{su}(2)}[\hbar]$, is an algebra of all polynomials in $X_{1},X_{2},X_{3}$ with relations $X_{i}X_{j}-X_{j}X_{i}-i\hbar\epsilon^{ijk}X_{k}~{}(i,j,k\in\{1,2,3\})$. Let $x_{a}~{}(1\leq a\leq 3)$ be commutative variables. $(x_{1},x_{2},x_{3})=(x,y,z)$ is identified with the coordinates of ${\mathbb{R}}^{3}$. The Lie-Poisson structure is defined by
$\{x_{a},x_{b}\}=i\epsilon^{abc}x_{c}.$ (5.1)
$A_{\mathfrak{su}(2)}$ is given by $\mathbb{C}[x]$ with this Poisson bracket. For arbitrary $f\in A_{\mathfrak{su}(2)}$ is given as
$f=f_{0}+f_{a}x_{a}+\frac{1}{2}f_{ab}x_{a}x_{b}+\cdots$
where $f_{a_{1}\cdots a_{i}}\in\mathbb{C}$ is completely symmetric with respect to $a_{1}\cdots a_{i}$. Let $V^{\mu}$ be a vector space ${\mathbb{C}}^{k}$. For example, we consider $V^{2}={\mathbb{C}}^{2}$, then the $q_{2}$ is given by a map from $A_{\mathfrak{su}(2)}$ to a matrix algebra ${\rm Mat}_{2}(\mathbb{C})$ is defined by
$q_{2}(f)$ $:=f_{0}{\bf 1}_{2}+f_{a}q_{2}(x_{a}),\quad\quad q_{2}(x_{a}):=\frac{\hbar}{2}\sigma^{a},$
where $\sigma^{a}$ is a Pauli matrix and ${\bf 1}_{k}$ is a $k\times k$ unit matrix.
In the case of $\dim V^{\mu}\geq 2$, the matrix regularization $q_{k}:A_{\mathfrak{su}(2)}\to{\rm Mat}_{k}(\mathbb{C})$ is defined by
$q_{k}(f)$ $:=f_{0}{\bf 1}_{k}+f_{a_{1}}q_{k}(x_{a_{1}})+\cdots+\frac{1}{k!}f_{a_{1}\cdots a_{k}}q_{k}(x_{a_{1}}\cdots x_{a_{k-1}})$
$q_{k}(x_{a_{1}}\cdots x_{a_{m}})$ $:=\frac{\hbar^{m}(k)}{m!}\sum_{\sigma\in Sym(m)}J_{a_{\sigma(1)}}\cdots J_{a_{\sigma(m)}}$
where $J_{a}$ are generators for the $k$-dimensional irreducible representation of $\mathfrak{su}(2)$ (the spin $2s+1=k$ representation). $q_{k}$ gives a map from a polynomial to a $k\times k$ matrix. $J_{a}$ satisfies
$[J_{a},J_{b}]$ $=i\epsilon^{abc}J_{c},\quad\quad[q_{k}(x_{a}),q_{k}(x_{b})]=i\hbar^{2}(k)\epsilon^{abc}J_{c}=i\hbar(k)\epsilon^{abc}q_{k}(x_{c}).$ (5.2)
Up to this point, we have not discussed the Casimir polynomial, so we have only been discussing the matrix regularization corresponding to the polynomial functions (Lie-Poisson algebra) defined on ${\mathbb{R}}^{3}$. If the series $\hbar(k)$ converging to 00 is given such that $\lambda^{2}$ in the equation (5.4) diverges, $q_{k}$ is a matrix regularization of $A_{\mathfrak{su}(2)}$ whose corresponding Lie-Poisson variety is ${\mathbb{R}}^{3}$.
From now on, we will consider the matrix regularization of the polynomial ring defined on the sphere by referring to the Casimir polynomial. Solving $\{x_{a},f(x)\}=0~{}(a=1,2,3)$ for a 2nd-degree homogeneous polynomial $f\in A_{\mathfrak{su}(2)}$, we obtain a solution as a quadratic Casimir polynomial
$f=\delta^{ab}x_{a}x_{b}.$ (5.3)
Then $q_{k}(f^{C}(x))=\hbar^{2}(k)\delta^{ab}J_{a}J_{b}$ is a Casimir invariant:
$\hbar^{2}(k)\delta^{ab}J_{a}J_{a}=\hbar^{2}(k)\frac{1}{4}(k^{2}-1){\bf 1}_{k}=:\lambda^{2}{\bf 1}_{k},$ (5.4)
where the eigenvalue $\lambda^{2}$ is a non-negative constant. We construct an ideal according to the method described in (4.21) and the following. So we chose the generators of the ideal $I(C)\subset A_{\mathfrak{su}(2)}$ as
$f^{C}(x):=\delta^{ab}x_{a}x_{b}-\lambda^{2},$ (5.5)
and
$I(C):=\left\{f^{C}(x)g(x)\in A_{\mathfrak{su}(2)}~{}|~{}g(x)\in A_{\mathfrak{su}(2)}~{}\right\}\subset A_{\mathfrak{su}(2)}.$ (5.6)
Then, $A_{\mathfrak{su}(2)}/I(C)$ is a set of polynomials on $S^{2}$ given by
$\delta^{ab}x_{a}x_{b}=\lambda^{2}.$ (5.7)
From (5.4) and Section 4.2, we find that when $\hbar(k)$ is chosen as
$\hbar(k)=\sqrt{\frac{4\lambda^{2}}{k^{2}-1}},$
then we can define a weak matrix regularization of $A_{\mathfrak{su}(2)}/I(C)$. To make a matrix regularization of $A_{\mathfrak{su}(2)}/I(C)$, we also build the other parts.
$I(C(X))$ :=$:=\left\{\sum_{i,j}a_{i}(X)f^{C}(X)b_{j}(X)\in\mathcal{U}_{\mathfrak{su}(2)}[\hbar]~{}|~{}a_{i}(X),b_{j}(X)\in\mathcal{U}_{\mathfrak{su}(2)}[\hbar]~{}\right\}.$ (5.8)
Here, $f^{C}(X)$ is a Casimir operator;
$f^{C}(X):=q_{U}(f^{C}(x))=\delta^{ab}X_{a}X_{b}-\lambda^{2}.$ (5.9)
The reduced Gröbner basis for this $I(C)$ is given by $G=\{\delta^{ab}X_{a}X_{b}-\lambda^{2}\}$. $\mathcal{U}_{\mathfrak{su}(2)}[\hbar]/I(C(X)),q_{U/I}$ and so on are defined by using these ideals and $G$. A linear function $\rho_{U/I,k}:\mathcal{U}_{\mathfrak{su}(2)}[\hbar]/I(C(X))\to gl(V^{k})=gl({\mathbb{C}}^{k})$ is defined as follows. For any monomial $X^{A}=X_{a_{1}}\cdots X_{a_{m}}~{}(m\in{\mathbb{N}})$ in $\mathcal{U}_{\mathfrak{su}(2)}[\hbar]$,
$\rho_{U/I,k}([X^{A}])=\hbar^{m}(k)J_{a_{1}}\cdots J_{a_{m}}.$ (5.10)
Finally the matrix regularization of $A_{\mathfrak{su}(2)}/I(C)$, is given as follows. For $f=h+r_{f,G}~{}(h\in I(C),r_{f,G}\notin I(C))$ with
$r_{f,G}=\sum_{A}c_{A}x^{A}=\sum_{\deg x^{A}\leq n_{k}}\!\!\!\!\!c_{A}x^{A}+\sum_{\deg x^{A}>n_{k}}\!\!\!\!\!c_{A}x^{A},$
the explicit calculation of $q_{A/I,k}:A_{\mathfrak{su}(2)}/I(C)\to gl(V^{k})$ is given as
$q_{A/I,k}([f(x)])=\sum_{m\leq n_{k}}\!c_{A}~{}\frac{\hbar^{m}(\mu)}{m!}\sum_{\sigma\in Sym(m)}J_{a_{\sigma(1)}}\cdots J_{a_{\sigma(m)}}.$ (5.11)
Thus, the matrix regularization of $A_{\mathfrak{su}(2)}/I(C)$, the fuzzy sphere, could be reconstructed by the method of this paper.
As an example, let us consider $f(x)=x^{3}\in A_{\mathfrak{su}(2)}$. By the graded lexicographic ordering, $r_{f,G}=x(-y^{2}-z^{2}+\lambda^{2})$. Then
$q_{A/I,k}([x^{3}])=R_{k}\left(-\frac{\hbar^{3}(k)}{3}(J_{1}J_{2}^{2}+J_{2}J_{1}J_{2}+J_{2}^{2}J_{1})-\frac{\hbar^{3}(k)}{3}(J_{1}J_{3}^{2}+J_{3}J_{1}J_{3}+J_{3}^{2}J_{1})+\hbar(k)\lambda^{2}J_{1}\right).$
If $n_{k}>3~{}(k>3)$, then $R_{k}$ is not different from a unit matrix in the above equation, i.e., $R_{k}(\cdots)=(\cdots)$, then the result is
$q_{A/I,k}([x^{3}])=\hbar^{3}(k)J_{1}^{3}+\frac{\hbar^{3}(k)}{3}J_{1}.$
For example, we consider $q_{3}:A_{\mathfrak{su}(2)}\to{\rm Mat}_{3}(\mathbb{C})$, generators $J_{a}$ for the $3$-dimensional irreducible representation of $\mathfrak{su}(2)$ given by
$J_{1}$ $=\begin{pmatrix}0&0&0\\ 0&0&i\\ 0&-i&0\end{pmatrix},\quad J_{2}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix},\quad J_{3}=\begin{pmatrix}0&i&0\\ -i&0&0\\ 0&0&0\end{pmatrix}.$
We construct the basis $E_{i}$ of ${\rm Mat}_{3}(\mathbb{C})$ according to the construction methods $1$ to $5$ in Subsection 3.2. First, the generators $E_{i}:=\hbar J_{i}~{}(i=1,2,3)$ and the unit matrix $E_{0}:=Id_{3}$ are chosen as basis elements. Next, we construct an independent element from the symmetrized product of $E_{i}$. That is,
$E_{4}:=$ $E_{1}^{2}=\frac{\hbar^{2}}{2}\begin{pmatrix}0&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\quad E_{5}:=E_{2}^{2}=\frac{\hbar^{2}}{2}\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&1\end{pmatrix},$
$E_{6}:=$ $\frac{1}{2}(E_{1}E_{2}+E_{2}E_{1})=\frac{\hbar^{2}}{2}\begin{pmatrix}0&-1&0\\ -1&0&0\\ 0&0&0\end{pmatrix},$
$E_{7}:=$ $\frac{1}{2}(E_{2}E_{3}+E_{3}E_{2})=\frac{\hbar^{2}}{2}\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&-1&0\end{pmatrix},$
$E_{8}:=$ $\frac{1}{2}(E_{1}E_{3}+E_{3}E_{1})=\frac{\hbar^{2}}{2}\begin{pmatrix}0&0&-1\\ 0&0&0\\ -1&0&0\end{pmatrix}.$
Since these $E_{i}~{}(i=0,\cdots,8)$ are independent of each other, we obtained a basis. Consider $f=x^{3}+x$ in $A_{\mathfrak{su}(2)}$. If $x>y>z$ with the graded lexicographic order as an ordering relation, then $r_{f,G}=\lambda^{2}x+x-xy^{2}-xz^{2}$ since the reduced Gröbner basis $G=\{z^{2}+y^{2}+x^{2}-\lambda^{2}\}$ from $f^{C}(x,y,z)=z^{2}+y^{2}+x^{2}-\lambda^{2}$ of (4.21) for $\mathfrak{su}(2)$. Note that $n_{3}=2$ and $\lambda^{2}=2\hbar^{2}(3)$,
$q_{A/I,3}([f])=\hbar(3)J_{1}=E_{1}.$
For another case, let $g=z^{3}+z$. Then $g=r_{g,G}=z^{3}+z$, and
$q_{A/I,3}([g])=\hbar(3)J_{3}=E_{3}.$
The commutator of these is obtained by
$[q_{A/I,3}([f]),q_{A/I,3}([g])]$ $=-i\hbar^{2}(3)J_{2}=-i\hbar(3)E_{2}.$
On the other hand, from $r_{\{f,g\},G}=i((-3\lambda^{2}-1)y-9\lambda^{2}yz^{2}+9y^{3}z^{2}+3y^{3}+9yz^{4})$, we obtain
$\hbar q_{A/I,3}(\{[f],[g]\})$ $=-i\hbar^{2}(3)J_{2}=-i\hbar(3)E_{2}.$
The difference between them is
$[q_{A/I,3}([f]),q_{A/I,3}([g])]-\hbar q_{A/I,3}(\{[f],[g]\})=0.$
5.2 $\mathfrak{su}(3)$; Fuzzy space
Let us consider $\mathfrak{su}(3)$ as $\mathfrak{g}$. For a typical example of representation of $\mathfrak{su}(3)$, generators $T_{i}$ of $\mathfrak{su}(3)$ Lie algebra are given by
$T_{i}=\frac{1}{2}\lambda_{i},$
where $\lambda_{i}$ are Gell-Mann matrices
$\lambda_{1}$ $=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\quad\lambda_{2}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},\quad\lambda_{3}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix},$
$\lambda_{4}$ $=\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{pmatrix},\quad\lambda_{5}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix},\quad\lambda_{6}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},$
$\lambda_{7}$ $=\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix},\quad\lambda_{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix}.$
From this basis, the structure constants are determined by
$f_{ab}^{c}=2~{}\mathrm{tr}[T_{a},T_{b}]T_{c}.$ (5.12)
In the following, this structure constant is fixed. Using this structure constant, the Lie-Poisson structure is given as
$\{x_{a},x_{b}\}=f^{c}_{ab}x_{c}$
for $x_{a}(a=1,2,\cdots,8)$. Solving $\{x_{a},f(x)\}=0~{}(a=1,\cdots,8)$ for a $2$nd-degree homogeneous polynomial $f\in A_{\mathfrak{su}(3)}$, we obtain a quadratic Casimir polynomial
$C^{2}(x)=\frac{1}{3}\delta^{ab}x_{a}x_{b}.$ (5.13)
For the other case, a cubic Casimir Polynomial is given as
$C^{3}(x)$ $=\frac{1}{18}(2\sqrt{3}x_{8}^{3}-6\sqrt{3}x_{1}^{2}x_{8}-6\sqrt{3}x_{2}^{2}x_{8}-6\sqrt{3}x_{3}^{2}x_{8}+3\sqrt{3}x_{4}^{2}x_{8}+3\sqrt{3}x_{5}^{2}x_{8}+3\sqrt{3}x_{6}^{2}x_{8}+3\sqrt{3}x_{7}^{2}x_{8}$
$\quad\quad-18x_{2}x_{5}x_{6}+18x_{2}x_{4}x_{7}-18x_{1}(x_{4}x_{6}+x_{5}x_{7})-9x_{3}\left(x_{4}^{2}+x_{5}^{2}-x_{6}^{2}-x_{7}^{2}\right)).$
Figure 1: Variety with $C^{3}(x)=1$. ($x_{2}=x_{3}=x_{4}=x_{5}=x_{7}=0$)
Thus, the possible spaces represented by the classical(commutative) limit are ${\mathbb{R}}^{8}$, $C^{2}(x)=const$ in ${\mathbb{R}}^{8}$, i.e. $S^{7}$, and $C^{3}(x)=const$ in ${\mathbb{R}}^{8}$.
i) Case of ${\mathbb{R}}^{8}$.
If the series $\hbar(\mu)$ converging to 00 is given such that $\lambda^{2}$ in the equation (5.13) diverges, $q_{\mu}$ is a matrix regularization of $A_{\mathfrak{su}(3)}$ whose corresponding Lie-Poisson variety is ${\mathbb{R}}^{8}$.
ii) Case of $S^{7}$ : (fixing each $\hbar(\mu)$ by $q_{\mu}(C^{2}(x)-\lambda^{2})=0$).
Consider $\dim V^{3}=3$ case as a simple example. $E_{i}=\hbar(3)T_{i}~{}(i=1,\cdots,8)$ and $E_{0}=Id_{3}$ yield a basis of ${\rm Mat}_{3}(\mathbb{C})$, so $n_{3}=1$. For example, $f=x_{1}+x_{1}^{2}$ in $A_{\mathfrak{su}(3)}$. In this case, the Gröbner basis is given by $\{C^{2}(x)-\lambda^{2}\}$, then
$r_{f,G}=x_{1}+3\lambda^{2}-x_{2}^{2}-x_{3}^{2}-x_{4}^{2}-x_{5}^{2}-x_{6}^{2}-x_{7}^{2}-x_{8}^{2}.$
Since $n_{3}=1$, we simply replace each $x_{i}$ with $\hbar(3)T_{i}$, then
$q_{A/I,3}([f])=E_{1}.$
Suppose $g(x)=x_{2}\in A_{\mathfrak{su}(3)}$. Since $g=r_{g,G}=x_{2}$, the commutator of them is given by
$[q_{A/I,3}([f]),q_{A/I,3}([g])]=$ $i\hbar(3)E_{3}.$
From the Lie-Poisson $\{[f],[g]\}=[ix_{3}+2ix_{1}x_{3}]$, the following is obtained:
$\hbar(3)q_{A/I,3}(\{[f],[g]\})=$ $\hbar(3)q_{A/I,3}([ix_{3}+2ix_{1}x_{3}])=i\hbar(3)E_{3}$
Therefore, in this case, $[q_{A/I,3}([f]),q_{A/I,3}([g])]=\hbar q_{A/I,3}(\{[f],[g]\})$. This result is expected from (4.6).
iii) Case of $C^{3}(x)=\lambda^{3}$ in ${\mathbb{R}}^{8}$: (fixing $\hbar(\mu)$ by $q_{\mu}(C^{3}(x)-\lambda^{3})=0$)
Consider the same case $\dim V^{\mu}=3$ as $S^{7}$ case. To distinguish it from Case ii) (case of $S^{7}$), we do not assign a specific number to $\mu$, and instead keep it as $\mu$ in the following discussion. Note also that the values of $\hbar(3)$ and $\hbar(\mu)$ are different. The basis is the same as in Case ii), i.e., $E_{i}=\hbar T_{i}~{}(i=1,\cdots,8)$ and $E_{0}=Id_{3}$ yield the basis. $n_{\mu}=1$ is also the same as in Case ii). The Gröbner basis is given by $\{C^{3}(x)-\lambda^{3}\}$. As an example, let us consider $f_{2}=x_{1}^{2}x_{3}x_{8}+x_{2}$. Then $r_{f_{2},G}$ is given by
$x_{2}-\sqrt{3}\lambda^{3}x_{3}-\sqrt{3}x_{1}x_{3}x_{4}x_{6}-\sqrt{3}x_{1}x_{3}x_{5}x_{7}-x_{2}^{2}x_{3}x_{8}+\sqrt{3}x_{2}x_{3}x_{4}x_{7}-\sqrt{3}x_{2}x_{3}x_{5}x_{6}-x_{3}^{3}x_{8}$
$+\frac{1}{2}\left(-\sqrt{3}x_{3}^{2}x_{4}^{2}-\sqrt{3}x_{3}^{2}x_{5}^{2}+\sqrt{3}x_{3}^{2}x_{6}^{2}+\sqrt{3}x_{3}^{2}x_{7}^{2}+x_{3}x_{4}^{2}x_{8}+x_{3}x_{5}^{2}x_{8}+x_{3}x_{6}^{2}x_{8}+x_{3}x_{7}^{2}x_{8}\right)+\frac{x_{3}x_{8}^{3}}{3}.$
Since $\lambda^{3}$ is in proportional to $\hbar^{3}(\mu)$, we get
$q_{A/I,\mu}([f_{2}])=\hbar(\mu)T_{2}=E_{2}.$
Let us consider $g_{2}=r_{g,G}=x_{4}$. The similar calculations as in the case of $S^{7}$ yield
$[q_{A/I,\mu}([f_{2}]),q_{A/I,\mu}([g_{2}])]$ $=\frac{i\hbar(\mu)}{2}E_{6}.$
The matrix regularization for Poisson brackets can also be performed straightforwardly, yielding the following,
$\hbar(\mu)q_{A/I,\mu}(\{[f_{2}],[g_{2}]\})$ $=[q_{A/I,\mu}([f_{2}]),q_{A/I,\mu}([g_{2}])].$
6 Summary
In this paper, the quantization of the Lie-Poisson algebra was carried out as a matrix regularization in a weak sense.
In Section 2, it was shown that the mass-deformed IKKT matrix model is equivalent to the matrix model whose solution is a basis of a semisimple Lie algebra. From this fact, a basis of every semisimple Lie algebra makes a classical solution of the mass-deformed IKKT matrix model. The precise statement of this claim is given in Theorem 2.2. Lie-Poisson algebras are expected as commutative limits of the algebras generated by these classical solutions. Matrix regularization connects the Lie-Poisson algebra and the algebra generated by a Lie algebra as a quantization. So, the matrix regularization of the Lie-Poisson algebras was studied. It is generalization of the fuzzy sphere. For a long time, the enveloping algebras of Lie algebras have been studied as a certain quantization of Lie-Poisson algebras. Giving a matrix representation of the enveloping algebra roughly corresponds to this matrix regularization. In this paper, we constructed a quantization by relaxing the standard conditions of matrix regularization. So we called it “weak matrix regularization” for the sake of distinction.
The process of constructing the weak matrix regularization for $A_{\mathfrak{g}}/I(C)$ is a little complicated, so it would be better to review the procedure here, where $\mathfrak{g}$ is a $d$-dimensional Lie algebra, and $A_{\mathfrak{g}}/I(C)$ is a Lie-Poisson algebra. The ideal $I(C)$ is not arbitrary, but is assumed to be made from $k$-th degree Casimir polynomials. These polynomials determine the geometry of $A_{\mathfrak{g}}/I(C)$. The following is a summary of the procedure for matrix regularization of $A_{\mathfrak{g}}/I(C)$.
1). At first we construct a matrix regularization with the trivial ideal $I(C)=\{0\}$, i.e., $A_{\mathfrak{g}}/I(C)=A_{\mathfrak{g}}$. This is given in Subsection 3.2.
We consider a representation of $\mathfrak{g}$ to $T_{\mu}\subset gl(V^{\mu})$ which is the algebra generated by the image of the representation. We set $n_{\mu}$ as a certain degree that determines the kernel of the quantization. The linear function $q_{\mu}:A_{\mathfrak{g}}\rightarrow T_{\mu}$ defined by $q_{\mu}(\sum_{k}f_{i_{1},\cdots,i_{k}}x^{i_{1}}\cdots x^{i_{k}})=\sum_{k}^{n_{\mu}}f_{i_{1},\cdots,i_{k}}e^{(\mu)}_{(i_{1},\cdots,i_{k})}$ is the matrix regularization for $A_{\mathfrak{g}}$.
2). We introduce a quantization map $q_{U}$ from $A_{\mathfrak{g}}$ to enveloping algebra $\mathcal{U}_{\mathfrak{g}}$ by $q_{U}(x_{\alpha_{1}}\cdots x_{\alpha_{m}})=X_{(\alpha_{1},\cdots,\alpha_{m})}$. This quantization is basically well known from long ago.
3). Next, we construct a quantization map $q_{U/I}:A_{\mathfrak{g}}/I(C)\to\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))$ for nontrivial $I(C)$, which is described in Section 4.1. $I(C)$ is not arbitrary, but is assumed to be made from $k$-th degree Casimir polynomials satisfying (3.27). We use $q_{\mu}$ in 1) to obtain the relation (3.27). Let $G$ be the reduced Gröbner basis of $I(C)$. For any $f(x)\in{\mathbb{C}}[x]$ $f(x)=r_{f}(x)+h_{f}(x)$ is uniquely determined by $G$, where $h_{f}(x)\in I(C)$ and $r_{f}(x)\notin I(C)$. Then we can define $q_{U/I}$ by $q_{U/I}([f(x)]):=[q_{U}(r_{f,G}(x))]$.
4). There exists an algebra homomorphism $\rho_{U/I,\mu}:\mathcal{U}_{\mathfrak{g}}[\hbar]/I(C(X))\to gl(V^{\mu})$. Using this $\rho_{U/I,\mu}$ and a projection operator $R_{\mu}:gl(V^{\mu})[\hbar(\mu)]\to gl(V^{\mu})[\hbar(\mu)]$ that restricts the degree of $\hbar$ to $n_{\mu}$ or less. Finally, we get the weak matrix regularization $A_{\mathfrak{g}}/I(C)\to gl(V^{\mu})$ by $q_{A/I,\mu}:=R_{\mu}\circ\rho_{U/I,\mu}\circ q_{U/I}$.
It is not only that the target space of the weak matrix regularization, $T_{\mu}$, and this Lie-Poisson algebra $A_{\mathfrak{g}}/I(C)$ are same structure as a Lie algebra. In the sense of Proposition 4.11 or Proposition 4.12, $T_{\mu}$ is “similar” to $A_{\mathfrak{g}}/I(C)$ in the limit as $\hbar(\mu)$ approaches zero. Here, the eigenvalues of the Casimir operators are fixed as (3.27) in the limit where $\hbar(\mu)$ approaches 00 and the dimension $\dim V^{\mu}$ approaches infinity. Therefore, it would be natural to think of the variety determined by the Casimir polynomials as the classical space realized in the limit where $\hbar$ is zero. However, the precision of this discussion of the classical(commutative) limit is a subject for future work.
To know how different the weak matrix regularization constructed in this paper is from the matrix regularization written in Appendix B using the operator norm, we still need to introduce the operator norm and examine each of the conditions in Appendix B. This is another future work that has not yet been started.
In addition, there is no established method for relating “obtaining an effective theory on a classical manifold as the low-energy limit of the IKKT matrix model” to “the space of the corresponding commutative limit in matrix regularization”. In fact, it was also seen in this paper that the manifold obtained in the commutative limit is not uniquely determined. Refining these discussions is also a future issue.
Acknowledgements
A.S. was supported by JSPS KAKENHI Grant Number 21K03258. The author also thanks the participants in the workshop “Discrete Approaches to the Dynamics of Fields and Space-Time” for their useful comments. We would like to thank A. Tsuchiya and J. Nishimura for important information of the IKKT matrix model.
Appendix A Definition of $\tilde{O}(z^{n})$
Since we have not defined a norm on algebras in this paper, the Landau symbol $O$ does not make sense. So, we define an order $\tilde{O}$ by $x\in\mathbb{R}$ using the absolute value of a complex number.
Definition A.1.
Let $\mathcal{V}$ be a vector space over $\mathbb{C}$. Let every $f_{i}$ be a complex valued continuous function such that
$\lim_{x\to 0}\left|\frac{f_{i}(xz)}{x^{n}}\right|<\infty,$
where $x\in\mathbb{R}$ and $z\in\mathbb{C}$. For $a_{i}\in\mathcal{V}$ which is independent of $z\in\mathbb{C}$, we denote the element described as $\sum_{i}f_{i}(z)a_{i}\in\mathcal{V}$ by $\tilde{O}(z^{n})$. If every $f_{i}$ satisfies
$\lim_{x\to 0}\frac{f_{i}(xz)}{x^{n}}=0,$
then we denote $\sum_{i}f_{i}(z)a_{i}\in\mathcal{V}$ by $\tilde{O}(z^{n+\epsilon})$.
Note that $z$ itself is not necessarily continuous. In this paper, the case where $\mathcal{V}$ is an algebra often appears, but we are applying the above definition by considering it as a vector space. For the purpose of this symbol, it is possible to replace $\tilde{O}(z^{n+1})$ with $\tilde{O}(z^{n})$, but $\tilde{O}(z^{n})$ must not be replaced with $\tilde{O}(z^{n+1})$, in the same way as for the usual ${O}(z)$.
This definition can also be extended to the case of a quotient space as follows.
Definition A.2.
Let $\mathcal{V}$ be an algebra over a commutative ring $R$. For some $h\in\mathcal{V},~{}[h]\in\mathcal{V}/\sim$, if there exist $h^{\prime}\in\mathcal{V}$ such that $[h]=[h^{\prime}]$ and $h^{\prime}$ is $\tilde{O}(\hbar^{n})$, then we say $[h]\in\mathcal{V}/\sim$ is $\tilde{O}(\hbar^{n})$.
Example A.3.
Let consider enveloping algebra $\mathcal{U}_{\mathfrak{g}}[\hbar]$ with $X_{i}X_{j}-X_{j}X_{i}\sim\hbar f_{ij}^{k}X_{k}$ introduced in Subsection 3.2. $[\hbar(X_{i}X_{j}-X_{j}X_{i})]=\hbar^{2}f_{ij}^{k}[X_{k}]$ is $\tilde{O}(\hbar^{2})$.
Example A.4.
Let us consider $\mathcal{U}_{\mathfrak{g}}(\hbar)/I(C(X))$ in Subsection 4.1. $[f(X)]\in\mathcal{U}_{\mathfrak{g}}(\hbar)/I(C(X))$ is said to be $\tilde{O}(\hbar^{n})$ if there exists a $h(X)\in\mathcal{U}_{\mathfrak{g}}(\hbar)$ such that $[h(X)]=[f(X)]$ and $h(X)$ is $\tilde{O}(\hbar^{n})$ in the sense of Example A.3.
The following fact is proved in [72].
Proposition A.5.
Let $t_{i}:A\rightarrow M_{i}$ be a quantization defined by Definition 3.1, and let $h_{ij}:M_{i}\rightarrow M_{j}$ be an $R$-algebra homomorphism. Then
$h_{ij}(\tilde{O}(\hbar^{1+\epsilon}(t_{i})))=\tilde{O}(\hbar^{1+\epsilon}(t_{i}))\in M_{j}.$
Appendix B Matrix regularization
In this section, let us review the definition of standard matrix regularization for a symplectic manifold $(M,~{}\omega)$ in order to compare it with the definition given in this paper. Matrix regularization [41] has evolved from the ideas of Berezin-Toeplitz quantization [26, 75], Fuzzy space [61], and so on. One mathematically sophisticated formulation is given in [71]. However, there is no unified common formulation. Here, we introduce one of the definitions of matrix regularization as described in [14], which is a widely known definition.
Definition B.1.
Let $N_{1},N_{2},\ldots$ be a strictly increasing sequence of positive integers and $\hbar$ be a real-valued strictly positive decreasing function such that $\lim_{N\to\infty}N\hbar(N)$ converges. Let $T_{k}$ be a linear map from $C^{\infty}(M)$ to $N_{k}\times N_{k}$ Hermitian matrices for $k=1,2,\ldots$. If the following conditions are satisfied, then we call the pair $(T_{k},~{}\hbar)$ a $C^{1}$-convergent matrix regularization of $(M,~{}\omega)$.
  1. 1.
    $\lim_{k\to\infty}\|T_{k}(f)\|<\infty$,
  2. 2.
    $\lim_{k\to\infty}\|T_{k}(fg)-T_{k}(f)T_{k}(g)\|=0$,
  3. 3.
    $\lim_{k\to\infty}\|\frac{1}{i\hbar(N_{k})}[T_{k}(f),T_{k}(g)]-T_{k}(\{f,g\})\|=0$,
  4. 4.
    $\lim_{k\to\infty}2\pi\hbar(N_{k}){\rm Tr}T_{k}(f)=\int_{M}f\omega$,
[…]
As in this definition, one of the main differences between the many definitions of matrix regularization and the definition used in this paper is the introduction of the operator norm. Definition B.1 also requires the recovery of homomorphism in the limit using the operator norm, whereas no such requirement is made in this paper. The paper also imposes no restrictions on integrals or traces. Overall, the definition of quantization in this paper is a less restrictive formulation. Therefore, the term “weak matrix regularization” is used to distinguish it. Originally, in the category-theoretic definition of the classical limit in [72], we developed a formulation capable of describing various quantization procedures in a unified framework. For this reason, the introduction of a norm is deliberately avoided, and the space is kept as structure-free as possible. This explains why fewer conditions are imposed compared to many definitions of matrix regularization.
Appendix C A Brief summary of Gröbner basis
The definitions of words and phrases related to the Gröbner basis and some of its properties are summarized in this appendix. (See for example [30]. )
Let $K$ be a field. In this paper we consider $K={\mathbb{C}}$. Let $K[x_{1},\dots,x_{n}]$ be a polynomial ring with some fixed order. For example, the graded lexicographic ordering for monomials $X^{\alpha}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}$ and $X^{\beta}=x_{1}^{\beta_{1}}x_{2}^{\beta_{2}}\cdots x_{n}^{\beta_{n}}$ ($\alpha=(\alpha_{1},\dots,\alpha_{n}),\beta=(\beta_{1},\dots,\beta_{n})$) is given by
$\alpha<\beta\xLeftrightarrow{def}\begin{cases}deg(X^{\alpha}):=\alpha_{1}+\cdots+\alpha_{n}<deg(X^{\beta}):=\beta_{1}+\cdots+\beta_{n}\\ \alpha_{1}=\beta_{1},\dots,\alpha_{i-1}=\beta_{i-1},\alpha_{i}<\beta_{i}~{}\mbox{ when}~{}\alpha_{1}+\cdots+\alpha_{n}=\beta_{1}+\cdots+\beta_{n}\end{cases}.$
We employ the graded lexicographic ordering as the fixed order in this paper.
Next, we introduce some terms to define the Gröbner basis.
Definition C.1.
Let $f=\sum_{\alpha}a_{\alpha}x^{\alpha}\quad\left(\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{n}),~{}x^{\alpha}=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}\right)$ is an element of $K[x_{1},\dots,x_{n}]$ with some fixed order.
  1. 1.
    We say that $\max\{\alpha\mid a_{\alpha}\neq 0\}$ is multi degree of $f$, and we denote it by $\text{multdeg}(f)$.
  2. 2.
    Leading Monomial : ${LM}(f):=x^{\text{multdeg}(f)}$ is called the leading monomial of $f$
  3. 3.
    Leading Coefficient : ${LC}(f):=a_{\text{multdeg}(f)}$ is called the leading coefficient of $f$.
  4. 4.
    Leading Term : ${LT}(f):=\text{LC}(f)\cdot{LM}(f)$ is called leading term of $f$.
In addition, we introduce following symbols for some subset $S\subset K[x_{1},\dots,x_{n}]$. ${LM}(S):=\{{LM}(f)\mid f\in S\}$. We denote the monomial ideal generated by $LM(S)$ by $\langle\text{LM}(S)\rangle$. We also use ${LT}(S):=\{{LT}(f)\mid f\in S\},$ and the ideal generated by $LT(S)$ is denoted by $\langle\text{LT}(S)\rangle$. Therefore, we find
$\langle\text{LM}(S)\rangle=\langle\text{LT}(S)\rangle.$
Definition C.2.
Let $I$ be an ideal of $K[x_{1},\dots,x_{n}]$. We say that $G=\{f_{1},\dots,f_{s}\}\subset I$ is a Gröbner basis if
$\langle\text{LM}(I)\rangle=\langle\text{LM}(G)\rangle=\langle\text{LM}(f_{1}),\dots,\text{LM}(f_{s})\rangle.$
In the following, we list some important properties about the Gröbner basis [33, 30].
Proposition C.3.
For any monomial ordering and any ideal $I$ that is not $\{0\}$, there exists a Gröbner basis of $K[x_{1},\ldots,x_{n}]$ that generates $I$.
Proposition C.4.
Let $G$ be a Gröbner basis of $K[x_{1},\ldots,x_{n}]$ that generates $I$. For any $f\in K[x_{1},\ldots,x_{n}]$,
  1. 1.
    There exists a polynomial $g\in I$, and a polynomial $r\notin I$ satisfying $f=g+r$, such that any monomial in $r$ is not divisible by any element of $LT(G)$. We call $r$ the remainder.
  2. 2.
    The above $g$ and $r$ are uniquely determined by $f$ and $G$.
It is necessary to be careful, as the result may change depending on the choice of the Gröbner basis $G$. However, it is possible to introduce a good Gröbner basis that fixes the arbitrariness of the choice of $G$.
Definition C.5.
If a Gröbner basis $G=\{f_{1},\dots,f_{s}\}\subset I$ satisfies the following two conditions i) ${LC}(f_{i})=1$ for all $i$ ,   ii) no term in $f_{i}\in G$ is divisible by $LM(f_{j})(i\neq j)$, then we say that $G$ is a reduced Gröbner basis.
Theorem C.6.
Let $I\neq\{0\}$ be a polynomial ideal. Then, for a given monomial ordering, there is a reduced Gröbner basis for $I$, and it is unique.
The fact that Theorem 4.1 holds for a reduced Gröbner basis defined above is used in this paper.